### v1

# PTR: A Pre-trained Language Model for Trajectory Recovery

Abstract.
Spatiotemporal trajectory data is vital for web-of-things services and is extensively collected and analyzed by web-based hardware and platforms. However, issues such as service interruptions and network instability often lead to sparsely recorded trajectories, resulting in a loss of detailed movement data. As a result, recovering these trajectories to restore missing information becomes essential. Despite progress, several challenges remain unresolved. First, the lack of large-scale dense trajectory data hampers the performance of existing deep learning methods, which rely heavily on abundant data for supervised training. Second, current methods struggle to generalize across sparse trajectories with varying sampling intervals, necessitating separate re-training for each interval and increasing computational costs. Third, external factors crucial for the recovery of missing points are not fully incorporated.
To address these challenges, we propose a framework called PTR. This framework mitigates the issue of limited dense trajectory data by leveraging the capabilities of pre-trained language models (PLMs). PTR incorporates an explicit trajectory prompt and is trained on datasets with multiple sampling intervals, enabling it to generalize effectively across different intervals in sparse trajectories. To capture external factors, we introduce an implicit trajectory prompt that models road conditions, providing richer information for recovering missing points. Additionally, we present a trajectory embedder that encodes trajectory points and transforms the embeddings of both observed and missing points into a format comprehensible to PLMs. Experimental results on two public trajectory datasets with three sampling intervals demonstrate the efficacy and scalability of PTR.
1. Introduction
A trajectory is represented as a sequence of timestamped locations that depict the movement of individuals and vehicles. With the advances in web-of-things services like ride-hailing and sharing bikes and web-based platforms like Google Maps and Mapbox in recent years, massive trajectory data is collected and analyzed. Such data plays a pivotal role in various applications, including urban planning (Wang et al., 2024), traffic management (Rehena and Janssen, 2018; Wang et al., 2021), and personalized location services (Zhao et al., 2020; Wang et al., 2020).
Collecting trajectory data involves the coordination of multiple devices (Chen et al., 2024). The process begins with continuously recorded GPS coordinates from devices like smartphones or IoT sensors at regular intervals. These data points are then sent to central servers or databases for storage and analysis. However, factors such as network transmission instability, device power-saving settings, and poor coverage in high-traffic or remote areas will result in the loss of trajectory points, making the collected data sparse. Such sparse trajectories failed to accurately describe movement behavior and route choices, thus limiting their effectiveness in trajectory-based applications. To mitigate this problem, it is essential to recover the missing points from sparse trajectories which improves the integrity of trajectory information. Such task is usually termed trajectory recovery.
Many works have been proposed to tackle trajectory recovery task, which can be divided into two categories: free space trajectory recovery (Chen et al., 2011, 2023a) and map-matched trajectory recovery (Wei et al., 2024; Chen et al., 2023b). Free space trajectory recovery aims to directly restore the coordinates of missing trajectory points but cannot ensure that the recovered trajectory is constrained to the road network. Typically, an additional map-matching process is required before these trajectories can be used in navigation-based applications. Map-matched trajectory recovery focuses on directly recovering trajectories on the road network. In this paper, we focus on the latter setting because it can be directly utilized by downstream tasks without the need for further processing. Despite the progress made by existing methods on map-matched trajectory recovery, they still face some challenges as listed below.
1) The availability of dense trajectories is often limited. Deep learning-based trajectory recovery models are inherently data-driven, requiring a substantial number of dense trajectories as labels for supervised training. However, due to the aforementioned factors, most collected and stored trajectories are sparse, resulting in a limited availability of large-scale dense trajectory data. As a result, the lack of sufficient dense trajectories for training makes existing deep learning models prone to overfitting and leads to diminished performance.
2) Existing models struggle to generalize to sparse trajectories with varying sampling intervals. Sparse trajectories often exhibit inconsistent sampling intervals due to network instability. However, most trajectory recovery models are trained on sparse trajectories with fixed sampling intervals, limiting their ability to generalize to trajectories with varying intervals. As a result, when confronted with trajectories that have different sampling intervals, these models require separate training for each interval to recover trajectories effectively, leading to significantly increased training costs.
Figure 1. An example of road network evolution process.
3) External factors related to the recovery of missing points are not incorporated. External factors, such as road conditions, provide valuable insights that can enhance trajectory recovery. For instance, in the scenario illustrated in Figure 1, given the observed points $p_{1}$ and $p_{4}$ and the road conditions at time $t_{1}$, it is more likely that the user took route $R_{1}$ rather than $R_{2}$. Additionally, based on the road conditions at times $t_{2}$ and $t_{3}$, as shown in Figures 1 (b) and (c), it becomes easier to infer that the user first accelerated and then decelerated between $p_{1}$ and $p_{4}$. However, existing methods only consider information from the known trajectory points, overlooking the significance of external factors.
To address the aforementioned challenges, we propose a Pre-trained Language Model for Trajectory Recovery (PTR). By leveraging a pre-trained language model (PLM) that has been trained on large-scale corpus datasets and fine-tuning it with a limited set of dense trajectories, PTR can effectively recover trajectories even when only scarce dense trajectory data is available. PTR incorporates two key components to generalize to sparse trajectories with varying sampling intervals: first, an explicit trajectory prompt, which uses natural language to describe both the trajectory recovery task and its characteristics. This activates the PLM’s ability to recover trajectories and recognize the trajectory’s sampling interval. Second, a joint training strategy, where PTR is trained on trajectory data with multiple sampling intervals, further enhancing its generalization capabilities. To account for external factors, we introduce an implicit trajectory prompt, which integrates rich contextual information, such as road conditions, to provide insights on missing points. Lastly, we propose a trajectory embedder that encodes both observed and missing trajectory points, transforming them into representations interpretable by PLMs. This allows the PLM to capture the spatiotemporal correlations within the trajectory.
Overall, the main contributions of our work can be summarized as follows:
  • We introduce PTR, a novel framework for trajectory recovery that utilizes PLMs to efficiently recover trajectories using a limited set of dense trajectories.
  • We propose explicit trajectory prompts that enable the PLM to accurately interpret the trajectory recovery task and determine the sampling intervals of the trajectories. Combined with a joint training strategy on trajectories with various sampling intervals, PTR generalizes effectively to sparse trajectories across different intervals.
  • We present an implicit trajectory prompt to model external factors influencing missing trajectory points and a trajectory embedder to transform the trajectory representation into a form that the PLM can process.
  • We conduct extensive experiments on two real-world trajectory datasets, each with three different sampling intervals. The results show that our proposed model achieves state-of-the-art performance in both accuracy and efficiency.
2. Related Work
2.1. Free-space Trajectory Recovery
Trajectory recovery in free space aims to directly restore the GPS coordinates of missing points. Early approaches relied on predefined rules to model mobile objects (Chen et al., 2016; Su et al., 2013). For example, Linear (Hoteit et al., 2014) assumes that vehicle movement follows uniform linear motion and employs linear interpolation for trajectory recovery. MPR (Chen et al., 2011) assumes vehicles travel along the most frequently used routes, utilizing historical trajectory data to determine route frequencies for recovery. Other studies (Banerjee et al., 2014; Zheng et al., 2012) use Markov models to capture spatial transitions between trajectory points. However, these methods are limited to modeling low-order transitions and struggle to capture the global spatial-temporal dependencies of trajectories.
Recent deep learning-based approaches have provided more effective solutions for trajectory recovery (Xu et al., 2017, 2022). DHTR (Wang et al., 2019) employs a GRU-based seq2seq model to analyze user route transitions, integrating Kalman filtering to improve recovery accuracy. AttnMove (Xia et al., 2021) uses attention mechanisms and LSTM to capture overall human movement preferences. PeriodicMove (Sun et al., 2021) models each user trajectory as a graph, employing graph neural networks to efficiently capture user preferences by leveraging historical route transitions. TERI (Chen et al., 2023a) introduces a two-stage framework based on transformers to recover trajectories with unequal spacing, while TrajBERT (Si et al., 2023) utilizes the BERT (Devlin et al., 2018) framework to recover implicit trajectories, focusing on coarser-grained outputs such as area or Wi-Fi IDs. Although these methods show promising results, they require an additional map-matching process to align the recovered trajectories with the road network before they can be applied in navigation systems. This two-stage process leads to error accumulation and introduces significant time overhead, as the map-matching step is time-consuming.
2.2. Map-matched Trajectory Recovery
Compared with the above methods, map-matched trajectory recovery incorporates the road network as input and aims to recover the trajectory directly onto the road. To this end, MTrajRec (Ren et al., 2021) first represents the trajectory point with a road segment and moving rate, utilizing the seq2seq-based multi-task framework to capture the spatiotemporal correlations within the trajectory. Following this idea, RNTrajRec (Chen et al., 2023b) further models the relationship between trajectories and road networks, designing a spatial-temporal transformer architecture to encode sparse trajectories. MM-STGED (Wei et al., 2024) models sparse trajectories from a graph perspective, considering both micro and macro semantic information of trajectories.
Although these methods demonstrate promising progress in trajectory recovery, their performance are still hindered by the poor availability of large-scale dense trajectories and limited in generalization to different sampling intervals discussed in Section 1.
3. Preliminaries
3.1. Definitions
Definition 3.1 (Trajectory).
A trajectory is defined as a series of timestamped locations, denoted as $\mathcal{T}=\langle p_{1},\cdots,p_{|\mathcal{T}|}\rangle$ where $p_{i}=(lat_{i},lng_{i},t_{i})$ represents the latitude and longitude coordinates of an object at the time $t_{i},i\in\{1,\cdots,|\mathcal{T}|\}$. $|\mathcal{T}|$ is the length of trajectory. The sampling interval of the trajectory $\mathcal{T}$ is $t_{i}-t_{i-1},i\in\{2,\cdots,|\mathcal{T}|\}$.
Definition 3.2 (Road Network).
A road network is defined as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes. Each node $v\in\mathcal{V}$ signifies an intersection that links various road segments, and each node possesses attributes of latitude and longitude. $\mathcal{E}$ represents the set of edges, where each edge $e\in\mathcal{E}$ denotes road segment that connecting two intersections. An edge can be characterized by its starting intersection $e.start\in\mathcal{V}$ and ending intersection $e.end\in\mathcal{V}$.
Definition 3.3 (Map-matched Trajectory).
By employing the map-matching algorithm, a trajectory $\mathcal{T}$ can be projected onto the road network and return the map-matched trajectory $\mathcal{T}_{m}$. As a result, each point of the map-matched trajectory aligns accurately with a particular road. The map-matched trajectory is represented as $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{|\mathcal{T}_{m}|}\rangle$, where each $q_{j}=(e_{j},r_{j},t_{j})$ signifies the vehicle’s location at time $t_{j},j\in\{1,\cdots,|\mathcal{T}_{m}|\}$. Here, $e\in\mathcal{V}$ is the matched road segment, and $r$ is the moving ratio, which quantifies the proportion of the distance traveled from the starting point of a road segment in relation to the total length of that road segment.
Figure 2. An illustration of map-matched trajectory, road segment $e$, and moving ratio $r$.
Example 3.4.
Figure 2 gives an example of a map-matched trajectory. For a sparse trajectory $\mathcal{T}_{s}=\langle p_{1},p_{4},p_{7}\rangle$ with the sampling interval of 90 seconds, its corresponding dense map-matched trajectory is represented by $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{7}\rangle$ with a sampling interval of 30 seconds. Take a matching point $q_{6}$ as an example. It falls on the road $e_{3}$, so its road segment is $e_{3}$. The distance of $q_{6}$ from the start point of the road is represented as $d_{6}$, while the total length of $e_{3}$ is denoted by $l_{3}$. Therefore, the moving ratio $r_{6}$ can be calculated as $r_{6}=\frac{d_{6}}{l_{3}}$.
Figure 3. The framework of PTR, consists of Trajectory Prompt, Trajectory Embedder, and PLM-based Trajectory Encoder.
3.2. Problem Statement
Map-matched Trajectory Recovery. Given a sparse trajectory $\mathcal{T}_{s}=\langle p_{1},\cdots,p_{|\mathcal{T}_{s}|}\rangle$ with a sampling interval of $\mu$. The goal of map-matched trajectory recovery is to reconstruct the dense map-matched trajectory $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{|\mathcal{T}_{m}|}\rangle$ with a sampling interval of $\epsilon$. Note that the sampling interval $\mu$ is larger than $\epsilon$.
4. Methodology
4.1. Overall Pipeline
In this paper, we present PTR, which fine-tunes a PLM pre-trained on a large-scale corpus to efficiently recover trajectories with limited dense data. The framework of PTR, shown in Figure 3, comprises three main components: trajectory prompts, a trajectory embedder, and a PLM-based trajectory encoder.
Trajectory prompts offer essential information from two key components. First, explicit trajectory prompts, expressed in natural language, outline the trajectory recovery task and key trajectory characteristics. These prompts activate the PLM’s ability to recover trajectories and identify varying sampling intervals in sparse trajectories. With joint training on trajectories across different intervals, PTR achieves robust generalization. Second, implicit trajectory prompts model road conditions, providing crucial external factors of missing points.
The trajectory embedder effectively enhances the PLM’s understanding of trajectories. It begins by normalizing sparse trajectories with varying sampling intervals $\mu$ into a uniform interval $\epsilon$. Next, each trajectory point, including both observed and missing points, is encoded into an embedding that captures both local and global semantics. Finally, the embedding is transformed into a format understood with the PLM.
In the PLM-based trajectory encoder, most of the encoder’s parameters are frozen to preserve the pre-trained knowledge, while the multi-head attention layer remains trainable, allowing the model to capture the spatiotemporal correlations within the trajectory. Finally, the output layer predicts the road segment $e$ and the moving ratio $r$ for the recovered map-matched trajectory at each step.
4.2. Trajectory Prompts
To enable PLMs to recover trajectories with different intervals and efficiently model external factors for missing trajectory points, we introduce explicit and implicit trajectory prompts.
4.2.1. Explicit Trajectory Prompt
The explicit trajectory prompt offers a structured textual description at both the task and trajectory levels, enabling PTR to adapt to varying sampling intervals. At the task level, it includes the following components:
  • ¡Task Part¿: Sparse trajectory recovery.
  • ¡Target Part¿: Output the road segment and moving ratio for each point in the trajectory.
[…]
At the trajectory level, the explicit trajectory prompt provides detailed characteristics of the trajectory, consisting of:
[…]
The ¡Time Part¿ provides the trajectory’s specific start and end times, helping the PLM understand the duration and potential time patterns, such as morning or evening peaks. The ¡Movement Part¿ supports the PLM in inferring the trajectory’s movement.
After obtaining the prompt for each part, we use the tokenizer and word token embedding in PLMs to convert text into embeddings. The embedding of ¡Task Part¿, ¡Target Part¿, ¡Content Part¿, ¡Time Part¿, and ¡Movement Part¿ are denoted by $\textbf{H}_{ta}$, $\textbf{H}_{tar}$, $\textbf{H}_{c}$, $\textbf{H}_{t}$, and $\textbf{H}_{m}$. These embeddings are then concatenated to form the overall explicit trajectory prompt embedding.
(1) $\textbf{H}^{e}=\textbf{H}_{ta}||\textbf{H}_{tar}||\textbf{H}_{c}||\textbf{H}_{t}||\textbf{H}_{m},$
where ||||| | is the concatenate operation.
4.2.2. Implicit Trajectory Prompt
We incorporate road conditions as external factors for missing trajectory points, as they provide insights into both the surrounding environment and the object’s movement, which are all helpful for trajectory recovery. For instance, vehicles typically slow down in congested areas and accelerate in less congested regions. Given the complexity and variability of real-world road conditions, which are difficult to express in natural language, we represent them as an implicit trajectory prompt.
To obtain the road conditions for each trajectory point, we first calculate the average road conditions across all areas and time intervals, then extract the relevant information for each point. Specifically, we divide the area of interest into a spatial grid of $I\times J$ cells and partition the time into $T$ slices. For each grid cell, we compute the total traffic, resulting in a regional flow matrix $\textbf{RC}\in\mathbb{R}^{I\times J\times T}$. To capture the spatiotemporal correlations among different regions, we apply a 2D convolution (2D CNN) along the spatial dimensions and a 1D CNN along the temporal dimension, yielding the road condition representation $\textbf{H}_{rc}\in\mathbb{R}^{I\times J\times T\times F}$, where $F$ is the number of features. Formally:
(2) $\mathbf{H}_{rc}=\text{Conv1d(Conv2d({RC}))}\in\mathbb{R}^{I\times J\times T\times F}$
Obviously, the road conditions of a trajectory point are indexed by its longitude, latitude, and timestamp. However, it is challenging for missing points as their geographic coordinates are unknown. To address this, we introduce a road condition diffusion strategy that approximates the missing conditions using neighboring traffic. The implementation details will be elaborated in Section 4.3.2.
Figure 4. An illustration of trajectory preprocessing layer.
4.3. Trajectory Embedder
To efficiently encode trajectory points and enable PLMs to better understand them, we introduce the trajectory embedder.
4.3.1. Trajectory Preprocessing Layer
For different sparse trajectories, their sampling intervals $\mu$ are not constant, such as 4 minutes, 2 minutes, or 1 minute. To convert these into a uniform sampling interval, we introduce a placeholder ‘[m]’ to mark missing trajectory points and generate a preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, where the length of $|\mathcal{T}^{\epsilon}|=\frac{p_{|\mathcal{T}_{s}|}.t-p_{1}.t}{\epsilon}$. Although the location of ‘[m]’ is unknown, its timestamp can still be calculated, as illustrated in Figure 4.
4.3.2. Trajectory Feature Extractor
To embed trajectory points, we propose the trajectory feature extractor. Given the preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, there are two cases for its trajectory point characteristics:
  • Case 1: the trajectory point $s\in\mathcal{T}^{\epsilon}$ is observed. i.e., $\exists\,k\in\{1,\cdots,|\mathcal{T}_{s}|\},p_{k}.t=s.t$.
  • Case 2: the location of trajectory point $s\in\mathcal{T}^{\epsilon}$ is missing. i.e., $s=[m]$, note that its time $s.t$ is known.
For case 1, we use Learnable Fourier Features (Tancik et al., 2020; Li et al., 2021) to encode the continuous latitude and longitude of $s$ into an $F$-dimensional vector using the feature mapping function $\Phi(x):\mathbb{R}\to\mathbb{R}^{F}$. Formally:
(3) $\Phi(x)=W_{\Phi}[\mathrm{cos}\text{ }xW_{r}||\mathrm{sin}\text{ }xW_{r}],$
where $W_{r}\in\mathbb{R}^{F/2}$ and $W_{\Phi}\in\mathbb{R}^{F\times F}$ represent the learnable parameters, and $x\in\{s.lat,s.lon\}$.
Using the feature mapping function $\Phi(\cdot)$, the relative information between points $x$ and $y$ can be captured through multiplication operations. This relative information helps infer crucial details like the distance between trajectory points, which is important for understanding movement. For instance, a larger distance between two points may indicate a higher likelihood of vehicle acceleration.
Additionally, considering that vehicles move within the road network, the relationship between a trajectory point and its surrounding road segments is critical. To measure the relationship between a trajectory point $s$ and a road segment $l$, we define a function $f(d_{s,l})$ based on their shortest distance:
(4) $f(d_{s,l})=\begin{cases}e^{-(\frac{d_{s,l}}{\kappa})^{2}}&\text{ if }d_{s,l}<\varphi_{dist},\\ 0&\text{ otherwise ,}\end{cases}$
where $d_{s,l}$ is the shortest distance between $s$ and road $l$, $\kappa$ is a hyperparameter, and $\varphi_{dist}$ is a distance threshold. Then, we obtain the road network representation $\mathbf{h}_{s}^{\text{road}}$ by:
(5) $\mathbf{h}_{s}^{\text{road}}=\frac{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})*\textbf{M}_{l}}{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})},$
where $\textbf{M}_{l}\in\mathbb{R}^{F}$ is the learnable embedding of road segment $l$, and $\mathcal{|V|}$ is the total number of road segments. Finally, the complete representation of a trajectory point $s$ is obtained as:
(6) $\mathbf{h}_{s}=W_{1}[(\Phi(s.lat)+\Phi(s.lng))||\mathbf{h}_{s}^{\text{road}}]+b_{1},$
where $W_{1}\in\mathbb{R}^{2F\times F}$ and $b_{1}\in\mathbb{R}^{F}$ is the learnable parameters.
For case 2, where the location of point $s$ is unknown, here we use road conditions to represent its features, as described in Section 4.2.2. Since the road conditions at a specific point are influenced by the surrounding conditions, which propagate along both temporal and spatial dimensions, we propose a road condition diffusion strategy to model the road conditions of the missing location.
First, for the missing point $s$, we identify the observed forward and backward trajectory points, $s_{f}$ and $s_{b}$, respectively. For instance, in Figure 4, suppose we have a missing point $s$ timestamped at 8:00:45, its observed forward point $s_{f}=p_{1}$ and the backward point $s_{b}=p_{2}$. Then, we retrieve the road conditions of $s_{f}$ and $s_{b}$ as follows:
(7) $\begin{split}\mathbf{h}_{rc}^{f}&=\mathbf{H}_{rc}[\pi_{lat}(s_{f}.lat),\pi_{lng}(s_{f}.lng),\pi_{t}(s_{f}.t)],\\ \mathbf{h}_{rc}^{b}&=\mathbf{H}_{rc}[\pi_{lat}(s_{b}.lat),\pi_{lng}(s_{b}.lng),\pi_{t}(s_{b}.t)],\end{split}$
where $\pi_{lat}$, $\pi_{lng}$, and $\pi_{t}$ are index functions mapping latitude, longitude, and time to their respective indices. Next, we calculate the time interval between $s$ and $s_{f}$ and $s_{b}$, denoted as $\triangle t_{f}=s.t-s_{f}.t$ and $\triangle t_{b}=s_{b}.t-s.t$. The road condition of point $s$ is then computed as:
(8) $\mathbf{h}_{rc}^{s}=\frac{e^{-\triangle t_{f}}\mathbf{h}_{rc}^{f}+e^{-\triangle t_{b}}\mathbf{h}_{rc}^{b}}{e^{-\triangle t_{f}}+e^{-\triangle t_{b}}}$
In addition, we encode the time intervals $\triangle t_{f}$ and $\triangle t_{b}$ to quantify the relative position of $s$. The final representation of $s$ is given by:
(9) $\begin{split}\mathbf{h}_{s}&=W_{2}[\textbf{m}\text{ }||\text{ FC}(\triangle t_{f}||\triangle t_{b})\text{ }||\text{ }\mathbf{h}_{rc}^{s}]+b_{2},\\ \text{FC}(\cdot)&:\mathbb{R}^{2}\to\mathbb{R}^{F},\end{split}$
where $\textbf{m}\in\mathbb{R}^{F}$ is a learnable vector to represent the location is missing, and FC()($\cdot$)( ⋅ ) is the fully connection layer. $W_{2}\in\mathbb{R}^{3F\times F}$ and $b_{2}\in\mathbb{R}^{F}$ are the learnable parameters. Finally, the overall trajectory representation is $\textbf{H}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times F}$.
4.3.3. Trajectory Feature Transformation
To enhance the model’s ability to comprehend trajectory features, we introduce $K$ reference tokens $\textbf{E}_{w}\in\mathbb{R}^{K\times F}$, inspired by (Lee et al., 2019; Guo et al., 2023), to bridge the connection between PLMs and the trajectory. These reference tokens are designed to capture the global semantics of the trajectory.
Given the trajectory embedding H, we first apply a 1D CNN to aggregate neighboring information and capture local movement patterns:
(10) H=Conv1d($\textbf{H})$ = Conv1d ( H )
Next, we compute self-attention between the trajectory embedding $\textbf{H}^{\prime}$ and the reference tokens $\textbf{E}_{w}$ to capture the global semantics, where $\textbf{H}^{\prime}$ acts as the query, and $\textbf{E}_{w}$ as both the key and value:
(11) $\textbf{H}^{\prime\prime}=\text{Attention}(\textbf{H}^{\prime},\textbf{E}_{w},\textbf{E}_{w})$
After transforming the embedding $\textbf{H}^{\prime\prime}$ in the token space, we concatenate it with the explicit trajectory prompt embedding $\textbf{H}^{e}$ to form the final trajectory representation:
(12) $\textbf{Z}=\textbf{H}^{e}||\textbf{H}^{\prime\prime}$
Then, we add the Transformer positional embedding $PE$ in each element of Z and feed it into PLMs encoder to encode trajectory.
4.4. PLM-based Trajectory Encoder
As shown in Figure 3, we use the stacked Transformer layers as the fundamental architecture of PLMs. To make PLMs fully adaptable to the trajectory recovery task, we employ the Low-Rank Adaptation (LoRA) algorithm (Hu et al., 2021). LoRA introduces additional parameters within the Transformer block, reducing the number of trainable parameters and computational complexity while enhancing the PLM’s performance on new tasks.
Specifically, in each self-attention block of the Transformer, we introduce additional weights $\Delta\textbf{W}_{q}\in\mathbb{R}^{F\times F}$, $\Delta\textbf{W}_{k}\in\mathbb{R}^{F\times F}$, and $\Delta\textbf{W}_{v}\in\mathbb{R}^{F\times F}$ for the query, key, and value matrices $\textbf{W}_{q}$, $\textbf{W}_{k}$, and $\textbf{W}_{v}$. To reduce the parameter count, the additional weights $\Delta\textbf{W}$ are decomposed into two low-rank matrices, $\textbf{B}\in\mathbb{R}^{F\times r}$ and $\textbf{C}\in\mathbb{R}^{r\times F}$, where $r\ll F$. The modified query, key, and value matrices in each self-attention block are updated as $\textbf{W}_{q}=\textbf{W}_{q}+\Delta\textbf{W}_{q}$, $\textbf{W}_{k}=\textbf{W}_{k}+\Delta\textbf{W}_{k}$, and $\textbf{W}_{v}=\textbf{W}_{v}+\Delta\textbf{W}_{v}$. Formally, the output of the $l$-th transformer layer is defined as:
(13) $\textbf{Z}^{l}=\text{LoRA(}\text{Transformer }(\textbf{Z}^{l-1}))$
After stacking $L$ Transformer layers, we discard the explicit trajectory prompt portion and obtain the output embeddings $\mathbf{Z}^{\prime}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times d}$.
For the $j$-th element of $\textbf{Z}^{\prime}$, we apply the softmax function to calculate the probability of each road segment and use the argmax function to return the predicted road segment $e_{j}$. We employ a MLP with a Sigmoid activation function to determine the moving ratio $r_{j}$.
4.5. Training
4.5.1. Loss Function
We use multi-task learning to simultaneously optimize road segment recovery and moving ratio recovery tasks. For the road segment recovery, we use the cross-entropy as the loss function:
(14) $\mathcal{L}_{e}=-\frac{1}{|\mathcal{T}^{\epsilon}|}\sum_{j=1}^{|\mathcal{T}^{\epsilon}|}\mathrm{log}(\hat{e}_{j}|\textbf{Z}^{\prime}_{j})$
For the moving ratio, we use the mean squared error as the loss function:
(15) $\mathcal{L}_{r}=\frac{1}{|\mathcal{T}^{\epsilon}|}\sum_{j=1}^{|\mathcal{T}^{\epsilon}|}|r_{j}-\hat{r}_{j}|^{2}$
The final loss function is $\mathcal{L}=\mathcal{L}_{e}+\lambda\mathcal{L}_{r}$, where $\lambda$ is a hyper-parameter to balance two tasks.
4.5.2. Training Strategy
To generalize PTR in recovering sparse trajectories with various sampling intervals, we propose the following training approach.
For the dense trajectory dataset $\mathbb{T}$, which is set as a 15-second sampling interval in our experiments. To create sparse trajectory datasets, we re-sample $\mathbb{T}$ at three different intervals: 1 minute, 2 minutes, and 4 minutes, resulting in three datasets: $\mathbb{T}_{1}$, $\mathbb{T}_{2}$, and $\mathbb{T}_{4}$. These are then combined to form a larger training dataset, $\mathbb{T}_{all}$.
(16) $\mathbb{T}_{all}=\mathbb{T}_{1}\cup\mathbb{T}_{2}\cup\mathbb{T}_{4}.$
By training PTR on $\mathbb{T}_{all}$, we achieve balanced performance across the different sampling intervals. Subsequently, we fine-tune the model on each individual sparse trajectory dataset to further improve trajectory recovery performance.
5. Experiments
In this section, we present comprehensive experiments to evaluate the effectiveness of PTR on two real-world trajectory datasets with three different sparse trajectory sampling intervals. These experiments are designed to answer the following key questions:
  • RQ1: How does PTR’s performance compare to other baseline methods across different sampling intervals of sparse trajectories?
  • RQ2: How well does PTR perform when trained with limited data?
  • RQ3: What is the contribution of each component in PTR to the overall model performance?
5.1. Datasets
We assess our model using two publicly available trajectory datasets from Chengdu, China, and Porto, Portugal. The Chengdu dataset [^1] https://outreach.didichuxing.com/ contains trajectory data collected in November 2016. The Porto dataset [^2] https://www.kaggle.com/competitions/pkdd-15-predict-taxi-service-trajectory-i/data includes trajectory from 442 taxis, collected from July 2013 to June 2014. Both datasets are standardized to a sampling interval of 15 seconds. We remove trajectories with travel times less than 5 minutes or exceeding 1 hour, as well as outlier trajectories. The road network data is obtained from the OpenStreetMap website [^3] http://www.openstreetmap.org/. We employ a map-matching algorithm (Newson and Krumm, 2009) to project the trajectory onto the road network and obtain the ground truth of the road segments and moving rates. The detailed overview of the dataset characteristics is provided in Table 4.
5.2. Baselines
To evaluate the effectiveness of our model, we compare PTR with ten baseline methods. These include five free space trajectory recovery models: HMM (Newson and Krumm, 2009) + ShortestPath, Linear (Hoteit et al., 2014) + HMM (Newson and Krumm, 2009), MPR (Chen et al., 2011) + HMM (Newson and Krumm, 2009), DHTR (Wang et al., 2019) + HMM (Newson and Krumm, 2009), and AttnMove (Xia et al., 2021) + Rule, as well as five map-matched trajectory recovery models: MTrajRec (Ren et al., 2021), T2vec (Li et al., 2018) + Decoder, T3s (Yang et al., 2021) + Decoder, RNTrajRec (Chen et al., 2023b), and MM-STGED (Wei et al., 2024). The details of each baseline are provided in Appendix B.
5.3. Evaluation Metrics
We employ five common metrics to assess the effectiveness of our model followed (Ren et al., 2021; Chen et al., 2023b; Wei et al., 2024). For road segment recovery, we utilize Accuracy (Acc), Recall, and Precision (Prec) metrics to evaluate. A higher result indicates a more precise road segment recovery. And employ Mean Absolute Error (MAE) and the Root Mean Square Error (RMSE) in assessing the recovered GPS coordinates. Lower values of these metrics indicate a higher level of accuracy in the recovery results. The details of each metric are provided in Appendix C.
Table 1. Performance comparison on two datasets with sampling intervals at 4 minutes, 2 minutes, and 1 minute, respectively.
The best results are highlighted in bold, while the underline indicates the second-best results.
Sampling Interval Methods Chengdu Porto
[…]
$\mu=4$ minutes $\to\epsilon=15$ seconds HMM + ShortestPath 26.85 28.64 29.55 939.3 1047.7 20.19 26.22 33.51 886.9 941.5
Linear + HMM 26.42 30.45 36.15 974.5 1145.4 32.23 36.09 49.80 489.3 637.3
MPR + HMM 36.93 38.62 44.53 821.9 914.1 32.22 38.67 48.07 534.0 700.2
DHTR + HMM 41.48 57.34 50.48 673.6 911.3 32.02 58.29 45.61 456.7 627.1
AttnMove + Rule 63.43 73.97 78.72 358.2 916.7 49.31 48.62 78.03 310.0 621.3
MTrajRec 65.79 75.14 78.42 315.1 904.4 52.36 60.39 77.28 266.1 590.1
T3s + Decoder 65.60 75.26 78.14 318.2 926.3 52.24 60.24 77.80 270.4 594.9
T2vec + Decoder 66.51 75.68 78.27 307.5 915.2 53.13 60.27 77.62 256.3 571.0
RNTrajRec 67.66 75.59 79.97 306.1 886.0 54.59 60.42 79.20 248.1 549.1
MM-STGED 70.64 76.04 81.63 266.2 829.7 57.30 59.48 80.21 222.8 510.4
PTR 74.12 79.63 86.46 262.8 483.0 57.61 59.15 82.19 200.9 376.9
PTR + FT 74.58 80.09 86.63 253.2 465.7 57.67 59.08 82.05 200.8 370.2
$\mu=2$ minutes $\to\epsilon=15$ seconds HMM + ShortestPath 33.85 47.89 48.31 754.1 826.2 27.30 40.05 46.00 647.3 747.0
Linear + HMM 43.78 45.35 48.77 816.9 1054.7 49.35 50.45 63.87 408.7 609.8
MPR + HMM 49.88 54.62 50.94 474.8 899.0 49.76 52.97 62.86 409.8 610.9
DHTR + HMM 47.17 60.16 51.73 662.0 912.2 43.76 65.25 52.10 385.3 578.0
AttnMove + Rule 71.98 77.42 80.67 291.1 764.8 61.39 60.90 82.98 213.4 468.5
MTrajRec 74.52 78.25 81.09 254.5 885.7 61.65 65.65 78.99 179.9 451.5
T3s + Decoder 74.62 78.95 81.79 242.2 857.5 61.75 65.53 79.14 181.3 461.2
T2vec + Decoder 75.69 78.86 81.68 231.6 783.6 62.24 65.77 78.97 173.9 438.0
RNTrajRec 75.80 79.35 81.86 218.5 757.0 63.39 65.84 79.25 171.3 433.9
MM-STGED 78.14 80.06 83.58 197.2 696.0 65.69 66.15 80.74 152.5 400.8
PTR 81.76 84.31 88.38 187.4 366.2 66.40 66.52 82.14 141.9 294.6
PTR + FT 82.29 84.59 88.72 181.2 349.0 66.72 66.87 82.38 141.7 293.5
$\mu=1$ minute $\to\epsilon=15$ seconds HMM + ShortestPath 35.92 67.92 60.16 529.7 638.2 34.75 48.46 48.67 527.2 659.3
Linear + HMM 68.59 65.66 66.67 707.4 1005.2 66.17 64.72 75.22 368.3 571.1
MPR + HMM 62.25 62.67 60.53 418.8 659.0 66.27 65.66 74.51 402.6 628.2
DHTR + HMM 51.09 63.40 50.14 584.7 750.4 52.98 69.60 57.17 420.4 625.8
AttnMove + Rule 79.60 81.55 82.75 194.4 752.6 72.07 69.59 80.41 156.6 360.7
MTrajRec 81.12 81.73 83.75 187.1 718.4 71.65 70.92 80.35 114.9 332.3
T3s + Decoder 80.90 82.78 83.15 187.1 713.0 71.78 71.61 80.16 110.1 328.0
T2vec + Decoder 81.69 81.90 83.88 185.6 714.1 71.86 71.10 80.48 114.2 334.7
RNTrajRec 81.88 82.09 84.84 177.9 702.5 72.31 71.88 80.57 110.3 325.7
MM-STGED 84.26 84.15 85.92 154.0 633.5 73.16 72.27 80.81 108.2 321.9
PTR 87.17 87.99 90.52 141.4 290.8 74.42 72.78 82.82 95.6 211.7
PTR + FT 88.15 88.82 91.04 138.6 289.1 75.35 73.87 83.28 95.2 220.3
5.4. Settings
We split the trajectory dataset into training, validation, and testing sets in a 7:2:1 ratio. Following previous work (Ren et al., 2021; Chen et al., 2023b; Wei et al., 2024), for dense map-matched trajectories, we set the sampling interval $\epsilon=15$ seconds. To generate sparse trajectories, for each dense trajectory, we retain the first and last points and create three sparse versions by setting the sampling intervals $\mu=4$ minutes, 2 minutes, and 1 minute, respectively.
During training, we integrate sparse trajectories with three different sampling intervals to train PTR. After training, we evaluate its performance in recovering trajectories across various sampling intervals to assess its generalizability. To further enhance performance, we fine-tune the trained model for each specific sampling interval, allowing for a more precise recovery.
We employ PyTorch (Paszke et al., 2019) framework to implement PTR, with a learning rate of 1e-4 and a batch size of 64. BERT (Devlin et al., 2018) is selected as the foundation model for PLMs, as its encoder-only structure is well-suited to the reconstruction task due to its effective use of bi-directional contextual information from the trajectory (Ren et al., 2024). For road condition extraction, the area of interest is divided into a $64\times 64$ grid, and the time dimension is partitioned into hourly intervals. The hidden state dimension is set to $F=512$. In Equation 4, we set $\kappa=15$ and $\varphi_{dist}=50$ meters. The model is trained for 50 epochs with early stopping, using a patience of 10 epochs. All experiments are conducted on NVIDIA RTX A4000 GPUs.
Table 2. Scalibility analysis on Chengdu dataset.
Sampling Interval Rate 20% 40% 60% 80% 100%
[…]
$\mu=4$ minutes $\to\epsilon=15$ seconds MTrajRec 60.73 1048.0 63.58 968.5 64.73 929.4 65.39 914.6 65.79 904.4
T3s + Decoder 60.49 1098.9 63.32 968.7 64.23 961.9 65.04 942.3 65.60 926.3
T2vec + Decoder 62.17 966.1 64.92 946.1 64.95 972.3 65.31 946.8 66.51 915.2
RNTrajRec 60.80 998.0 62.85 931.3 64.19 909.1 65.46 835.0 67.66 886.0
MM-STGED 64.61 935.0 67.76 881.8 68.53 853.7 69.95 825.3 70.64 829.7
PTR 68.30 585.3 71.57 515.6 72.74 512.5 73.56 488.3 74.12 483.0
$\mu=2$ minutes $\to\epsilon=15$ seconds MTrajRec 69.85 919.8 72.58 907.4 73.97 902.3 74.32 891.3 74.52 885.7
T3s + Decoder 68.86 909.5 72.29 897.2 73.63 883.8 74.53 868.8 74.62 857.5
T2vec + Decoder 69.47 905.4 72.77 858.3 73.80 855.8 74.89 831.3 75.69 783.6
RNTrajRec 68.05 991.0 70.20 856.9 71.17 834.4 72.62 795.3 75.80 757.0
MM-STGED 71.65 865.0 75.32 773.0 75.49 752.1 76.94 747.7 78.14 696.0
PTR 76.29 452.8 79.37 411.5 80.37 395.4 81.18 377.4 81.76 366.2
$\mu=1$ minute $\to\epsilon=15$ seconds MTrajRec 75.39 937.5 78.53 835.1 80.01 794.3 80.93 725.7 81.12 718.4
T3s + Decoder 76.49 917.1 79.08 824.9 80.65 767.8 80.82 716.5 80.90 713.0
T2vec + Decoder 75.89 845.4 79.09 746.8 79.41 752.3 81.21 742.2 81.69 714.1
RNTrajRec 75.65 846.2 79.48 784.8 79.61 769.3 81.74 742.5 81.88 702.5
MM-STGED 76.02 857.7 79.86 734.7 82.25 676.4 83.44 663.0 84.26 633.5
PTR 81.62 397.5 84.59 352.0 85.75 328.7 86.73 305.4 87.17 290.8
5.5. Experimental Results (RQ1)
Table 1 reports a comparison of the results between our model and the baselines across various sampling intervals of sparse trajectories on both the Chengdu and Porto datasets. As the sampling interval increases, the map-matched trajectory recovery task is more difficult. Due to the more complex road network structure in Porto compared to Chengdu, with Porto having 100.9 roads / $km^{2}$ and Chengdu having 38.1 roads / $km^{2}$, the accuracy of road segment recovery is lesser in Porto than in Chengdu.
Compared to other baselines, our model achieves superior performance across all metrics, with an average improvement of 16.51% on the Chengdu dataset and 9.35% on the Porto dataset. Notably, it shows substantial gains in RMSE metrics, with an average reduction of 351.8 meters in Chengdu and 119.2 meters in Porto. This indicates PTR generalizes effectively, enabling it to accurately recover trajectories with varying sampling intervals. By fine-tuning at the specific sampling interval, PTR + FT, our model further improves Acc by 1.72% and 0.32%. These outstanding results can be attributed to the strong generalization capability of the PLM in processing sparse trajectories with varying sampling intervals. Additionally, the trajectory embedder plays a pivotal role in connecting trajectory data with the PLM, efficiently capturing spatial-temporal correlations and converting trajectory embeddings into a format that the PLM can effectively process and understand.
We also analyze the convergence rate of PTR across different sampling intervals, as shown in Figure 5 and  6. It demonstrates rapid convergence within just a few epochs on both the Chengdu and Porto datasets. While fine-tuning enhances performance further, the model stabilizes by the 3th epoch, indicating that PTR can deliver satisfactory results within a short training period.
5.6. Scalibility Analysis (RQ2)
To evaluate the effectiveness of PTR in trajectory recovery when dense trajectory data is limited, we conduct scalability experiments on the Chengdu dataset. Specifically, we train the model using subsets of the training set at 20%, 40%, 60%, 80%, and 100%, and evaluate performance on the test set. The results are presented in Table 2.
As the size of the training set increases, we observe performance improvements across all baseline models. Notably, with only 20% of the training data, PTR already surpasses most baselines, trailing only MM-STGED. With 40% of the training data, PTR outperforms all other models, demonstrating its strong generalization ability. This advantage underscores the practicality of PTR in real-world scenarios, where obtaining large-scale dense trajectory data is often challenging.
Table 3. Ablation study on two datasets with sampling intervals at 2 minutes.
[…]
Chengdu w/o explicit trajectory prompt 81.04 84.08 88.07 205.8 394.2
w/o implicit trajectory prompt 81.43 84.28 88.14 191.4 371.1
w/o trajectory prompt 80.70 83.61 87.73 225.9 407.4
w/o road network 81.28 84.22 88.04 194.5 375.6
w/o reference tokens 80.83 83.36 86.68 236.6 437.2
PTR 81.76 84.31 88.38 187.4 366.2
Porto w/o explicit trajectory prompt 66.13 66.19 81.98 157.9 314.6
w/o implicit trajectory prompt 66.18 66.22 82.01 149.6 298.1
w/o trajectory prompt 65.52 65.89 81.51 171.0 339.4
w/o road network 66.11 66.27 81.90 150.7 302.8
w/o reference tokens 65.89 65.07 80.57 165.4 327.7
PTR 66.40 66.52 82.14 141.9 294.6
5.7. Ablation Study (RQ3)
To evaluate the effects of various components in PTR, we conduct ablation experiments on the Chengdu and Porto datasets with sampling intervals of 2 minutes. We compare PTR with the following five variants.
  • w/o explicit trajectory prompt: We remove the explicit trajectory prompt.
  • w/o implicit trajectory prompt: We use the ‘[MASK]’ token in the BERT to represent the missing location.
  • w/o trajectory prompt: We remove both the explicit and implicit trajectory prompt, and the missing points are represented by the ‘[MASK]’ token in BERT.
  • w/o road network: We only use the Learnable Fourier Feature for encoding the observed trajectory points while removing the road network feature.
  • w/o reference tokens: We remove the trajectory feature transformation layer in the trajectory embedder module.
As shown in Table 3, each component of PTR is essential. The explicit trajectory prompt provides task and trajectory-related textual information, enabling PLMs to effectively understand the trajectory recovery tasks. Omitting this component leads to a performance decrease. When the implicit trajectory prompt is removed, many missing locations are replaced by the ‘[MASK]’ token in BERT, limiting the model’s ability to capture external factors for missing points, which results in suboptimal performance.
For the trajectory point embedder, excluding road network for observed points causes the model to rely solely on latitude and longitude, neglecting the relationship between trajectories and the road network, which demonstrates that road network is crucial for accurately modeling trajectory information. Additionally, removing reference tokens results in a performance decline, as PLMs are unable to directly interpret trajectory data.
6. Conclusion
In this paper, we propose PTR, a novel framework for trajectory recovery leveraging PLM. PTR efficiently handles trajectory recovery even with limited dense trajectory data, demonstrating strong scalability. The model is trained on datasets with various sparse trajectories, enabling it to effectively generalize to different sampling intervals. To model external factors, we introduce an implicit trajectory prompt to model road conditions, which provides rich information for missing trajectory points. We also develop a trajectory embedder to capture the spatial-temporal dependencies of sparse trajectories and convert them into a format interpretable by PLMs. Experimental results on two datasets with three different sampling intervals validate the efficacy and scalability of the proposed model.
Appendix A Dataset
Table 4. Dataset Description.
Types Chengdu Porto
Sampling interval 15s 15s
# Trajectory 118,354 322,079
# Road segment 2504 2224
Latitude range 30.655 $\sim$ 30.727 41.142 $\sim$ 41.174
Longitude range 104.043 $\sim$ 104.129 -8.652 $\sim$ -8.578
Length / Width 8.22 km / 8.00 km 6.19 km / 3.56 km
Appendix B Baseline Setting
We choose the following ten methods as baselines, including five free space trajectory recovery and five map-matched trajectory recovery.
B.1. Free-space Trajectory Recovery
Free space trajectory recovery first recovers trajectory points and then projects the trajectory onto the road network.
  • HMM (Newson and Krumm, 2009) + ShortestPath first projects the sparse trajectory on the road network based on the hidden markov model (HMM), and then calculate the shortest path.
  • Linear (Hoteit et al., 2014) + HMM (Newson and Krumm, 2009) linearly interpolates missing trajectory points and then implements HMM to perform the map matching process.
  • MPR (Chen et al., 2011) + HMM (Newson and Krumm, 2009) first divides the area of interest into grids to identify frequently traveled routes between sparse trajectory points. Then, it assumes that vehicle movement maintains a constant speed for trajectory recovery. Subsequently, it employs HMM to project the trajectory onto the road network.
  • DHTR (Wang et al., 2019) + HMM (Newson and Krumm, 2009) incorporates a sequence-to-sequence framework and Kalman filtering to recover trajectory points. Subsequently, it applies an HMM to yield a trajectory that is constrained by the map.
  • AttnMove (Xia et al., 2021) + Rule uses attention to predict the missing road segments and uses the central location as the moving rate.
B.2. Map-matched Trajectory Recovery
Map-matched trajectory recovery can directly recover the trajectory on the road network.
  • MTrajRec (Ren et al., 2021) utilizes a sequence-to-sequence framework with Gated Recurrent Units (GRU) as the key component for trajectory recovery. It optimizes road segment and moving rate prediction through multi-task learning.
  • T2vec (Li et al., 2018) + Decoder is a deep learning model for trajectory similarity learning. We use its encoder to embed the sparse trajectory and use the decoder of the MTrajRec to recover the trajectory in an autoregressive manner.
  • T3s (Yang et al., 2021) + Decoder uses LSTM and attention mechanisms to encode sparse trajectory data effectively. Then, the MTrajRec decoder is utilized to recover the trajectory.
  • RNTrajRec (Chen et al., 2023b) leverages the Transformer to capture the spatial-temporal correlation of the sparse trajectories. It also takes into account the relation between the trajectory and the road network.
  • MM-STGED (Wei et al., 2024) models sparse trajectories from a graph perspective and recovers trajectories by capturing micro and macro semantic information.
Appendix C Evaluation Metrics
We adopt five widely used metrics to evaluate the effectiveness of our model, following previous works (Wei et al., 2024; Chen et al., 2023b; Ren et al., 2021). For road segment recovery, we use Accuracy (Acc), Recall, and Precision (Prec) to assess the alignment between the true road segments $\mathcal{E}_{p}=\{e_{1},\cdots,e_{m}\}$ and the predicted road segments ^$\hat{\mathcal{E}}_{p}=\{\hat{e}_{1},\cdots,\hat{e}_{m}\}$. A higher value in these metrics indicates a more accurate road segment recovery. The metrics are formally defined as follows:
(17) $\begin{split}\textbf{Acc}&=\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{e_{i}=\hat{e}_{i}\}\times 100\%,\\ \textbf{Recall}&=\frac{|\mathcal{E}_{p}\cap\hat{\mathcal{E}}_{p}|}{|\mathcal{E}_{p}|}\times 100\%,\\ \textbf{Prec}&=\frac{|\mathcal{E}_{p}\cap\hat{\mathcal{E}}_{p}|}{|\hat{\mathcal{E}}_{p}|}\times 100\%,\end{split}$
where $\mathbbm{1}\{\cdot\}$ is the indicator function, where $e_{i}=\hat{e}_{i},\mathbbm{1}\{e_{i}=\hat{e}_{i}\}=1$, elsewise $\mathbbm{1}\{e_{i}=\hat{e}_{i}\}=0$.
To evaluate the recovered GPS coordinates, we employ the Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) to quantify the distance error between the true trajectory $\mathcal{T}_{m}={q_{1},\cdots,q_{|\mathcal{T}_{m}|}}$ and the predicted trajectory $\hat{\mathcal{T}}_{m}={\hat{q}_{1},\cdots,\hat{q}_{|\hat{\mathcal{T}}_{m}|}}$. The formulas for MAE and RMSE are given as follows:The formulas of MAE and RMSE are as follows:
(18) $\begin{split}\textbf{MAE}&=\frac{1}{|\mathcal{T}_{m}|}\sum_{i=1}^{|\mathcal{T}_{m}|}|\mathrm{RN\_dist}(q_{i},\hat{q}_{i})|,\\ \textbf{RMSE}&=\sqrt{\frac{1}{|\mathcal{T}_{m}|}\sum_{i=1}^{|\mathcal{T}_{m}|}|\mathrm{RN\_dist}(q_{i},\hat{q}_{i})|^{2}}\end{split}$
Here, following (Ren et al., 2021; Chen et al., 2023b; Wei et al., 2024), $\mathrm{RN\_dist}(q,\hat{q})$ signifies the shortest distance along the road network between the trajectory points $q$ and $\hat{q}$. Both MAE and RMSE are denoted in meters. Lower values of these metrics indicate a higher level of accuracy in the recovery results.
Appendix D Appendix Experiments
D.1. Covergence Rate
Figure 5. Convergence rate of PTR and fine-tuning for three sampling intervals on Chengdu.
Figure 6. Convergence rate of PTR and fine-tuning for three sampling intervals on Porto.
We evaluate the convergence speed of our model on the Porto and Chengdu datasets. As illustrated in Figure 5 and 6, the model reaches near-optimal performance by the 10th epoch. The trained model demonstrated robust accuracy across three sampling intervals (PTR: $u=1$ minute, PTR: $u=2$ minutes, and PTR: $u=4$ minutes), confirming that PTR can generalize effectively to sparse trajectories with varying sampling intervals. Furthermore, after fine-tuning at each sampling interval, the performance shows minimal change in performance by the 3rd epoch, further proving its fast convergence speed.
Table 5. Hyperparameter range and optimal value.
Parameter Range
The number of reference tokens $K$ 128, 256, 512, 1024
LoRA rank $r$ 4, 8, 16, 64
The loss function weight $\lambda$ 0.1, 1, 10, 100
Figure 7. Hyperparameter analysis of the number of reference tokens $K$ of trajectory feature transformation.
Figure 8. Hyperparameter analysis of the rank $r$ of LoRA.
Figure 9. Hyperparameter analysis of the weight $\lambda$ in the loss function.
D.2. Hyperparameter Study
To explore the impact of hyperparameters on model performance, we conduct hyperparameter analysis using the Chengdu and Porto datasets with the sparse trajectory sampling interval of 2 minutes. We consider three important hyperparameters, including the number of reference tokens $K$, the rank $r$ of LoRA, and the loss function weight $\lambda$. Details regarding the range of selection and the optimal values for these hyperparameters are presented in Table 5.
D.2.1. The Number of Reference Tokens $K$
We set the parameter $K$ from the set $\{128,256,512,1024\}$ to explore its impact on trajectory recovery. The experimental results are shown in Figure 7. As $K$ increases, the effectiveness of trajectory recovery also improves. This suggests that having a larger token space aids in accurately representing trajectory features. When $K$ is larger than 512, the performance improvement of PTR becomes marginal, while the computational cost will increase. To balance the performance and efficiency of the model, we set $K$ to 512.
D.2.2. The Rank $r$ of LoRA
As shown in Figure 8, we set the range of $r$ to 4, 8, 16, and 64 to explore its impact. It is observed that the model performs well when $r=4$. As $r$ increases, the accuracy will continue to increase, but it is not obvious. This suggests that PLMs encapsulate substantial domain expertise through training on extensive corpora, rendering them adaptable to trajectory recovery tasks with minor adjustments. Yet as $r$ increases, the number of parameters also increases, making the model training require more memory. Therefore, we set $r=8$ to balance the performance and resources.
D.2.3. The Loss Weight $\lambda$
As shown in Figure 9, as the loss function weight $\lambda$ increases, the model performance first improves and then decreases. This is because a smaller $\lambda$ will make the model focus on road segment recovery, while a larger $\lambda$ will focus on moving rate recovery. To balance these two tasks, we set the value of $\lambda$ to 10.
Figure 10. Case study on the Chengdu dataset. Red points represent the truth trajectory points and blue points represent the recovered trajectory points.
D.3. Case Study
We conduct a case study on the Chengdu dataset to visualize the trajectory recovery performance with various baselines. As shown in Figure 10, we draw the truth and recovered trajectory, where red points represent the truth trajectory points and blue points indicate the recovered trajectory points. We find that the recovered trajectory aligns well with the road network, demonstrating the effectiveness of using road segment and moving rate to represent trajectory point. To facilitate a more intuitive comparison of recovery performance among different models, we focus on two distinct regions, labeled as A and B, for visual analysis. In region A, characterized by a relatively simple road network structure, all models can accurately recover road segments. Among them, PTR maps trajectory points better by leveraging PLMs. In contrast, region B exhibits a more intricate road network with multiple accessible routes. RNTrajRec captures spatial-temporal correlations of trajectories that are not adequate and recovers incorrect road segments. PTR not only excels in road segment recovery but also in accurately matching the actual trajectory points. This success can be attributed to its ability to model missing trajectory points through the implicit trajectory prompt, thereby introducing more valuable information and improving performance.
D.4. Computational Cost
We use an NVIDIA RTX A4000 card to conduct cost analysis on the Chengdu and Porto datasets with a 2-minute sampling interval by considering model size, training time, and inference time. As shown in Table 6, compared to the end-to-end deep learning model, PTR has the largest number of parameters due to its utilization of a pre-trained language model based on the BERT-small framework. Despite having a greater number of parameters, PTR exhibits quicker training and testing times as it can output results for all trajectory points at once, and the majority of parameters are frozen, unlike other models that necessitate autoregressive generation. It significantly accelerates both model training and inference processes.
Table 6. Computation Cost.
Dataset Chengdu / Porto
Methods Model size Train time Inference time
(MBytes) (min/epoch) (min)
DHTR 19 / 18 11.30 / 42.55 1.13 / 2.35
MTrajRec 49 / 44 37.19 / 142.12 12.57 / 28.20
RNTrajRec 112 / 109 72.30 / 153.29 29.14 / 43.58
MM-STGED 34 / 31 41.27 / 138.33 16.27 / 30.58
PTR 176 / 175 24.18 / 54.14 2.56 / 10.38

### v2

@@
-PTR: A Pre-trained Language Model for Trajectory Recovery
+PLMTrajRec: A Scalable and Generalizable Trajectory Recovery Method with Pre-trained Language Models
 Abstract.
-Spatiotemporal trajectory data is vital for web-of-things services and is extensively collected and analyzed by web-based hardware and platforms.
-However, issues such as service interruptions and network instability often lead to sparsely recorded trajectories, resulting in a loss of detailed movement data.
-As a result, recovering these trajectories to restore missing information becomes essential.
-Despite progress, several challenges remain unresolved.
-First, the lack of large-scale dense trajectory data hampers the performance of existing deep learning methods, which rely heavily on abundant data for supervised training.
-Second, current methods struggle to generalize across sparse trajectories with varying sampling intervals, necessitating separate re-training for each interval and increasing computational costs.
-Third, external factors crucial for the recovery of missing points are not fully incorporated.
-To address these challenges, we propose a framework called PTR.
-This framework mitigates the issue of limited dense trajectory data by leveraging the capabilities of pre-trained language models (PLMs).
-PTR incorporates an explicit trajectory prompt and is trained on datasets with multiple sampling intervals, enabling it to generalize effectively across different intervals in sparse trajectories.
-To capture external factors, we introduce an implicit trajectory prompt that models road conditions, providing richer information for recovering missing points.
-Additionally, we present a trajectory embedder that encodes trajectory points and transforms the embeddings of both observed and missing points into a format comprehensible to PLMs.
-Experimental results on two public trajectory datasets with three sampling intervals demonstrate the efficacy and scalability of PTR.
+Spatiotemporal trajectory data is crucial for various applications.
+However, issues such as device malfunctions and network instability often cause sparse trajectories, leading to lost detailed movement information.
+Recovering the missing points in sparse trajectories to restore the detailed information is thus essential.
+Despite recent progress, several challenges remain.
+First, the lack of large-scale dense trajectory data makes it difficult to train a trajectory recovery model from scratch.
+Second, the varying spatiotemporal correlations in sparse trajectories make it hard to generalize recovery across different sampling intervals.
+Third, the lack of location information complicates the extraction of road conditions for missing points.
+To address these challenges, we propose a novel trajectory recovery model called PLMTrajRec.
+It leverages the scalability of a pre-trained language model (PLM) and can be fine-tuned with only a limited set of dense trajectories.
+To handle different sampling intervals in sparse trajectories, we first convert each trajectory’s sampling interval and movement features into natural language representations, allowing the PLM to recognize its interval.
+We then introduce a trajectory encoder to unify trajectories of varying intervals into a single interval and capture their spatiotemporal relationships.
+To obtain road conditions for missing points, we propose an area flow-guided implicit trajectory prompt, which models road conditions by collecting traffic flows in each region.
+We also introduce a road condition passing mechanism that uses observed points’ road conditions to infer those of the missing points.
+Experiments on two public trajectory datasets with three sampling intervals each demonstrate the effectiveness, scalability, and generalization ability of PLMTrajRec.
+Trajectory recovery, Pre-trained language model, Road network
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: KDD 2025; August 03–07, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/XX/XX
 1. Introduction
-A trajectory is represented as a sequence of timestamped locations that depict the movement of individuals and vehicles.
-With the advances in web-of-things services like ride-hailing and sharing bikes and web-based platforms like Google Maps and Mapbox in recent years, massive trajectory data is collected and analyzed.
-Such data plays a pivotal role in various applications, including urban planning (Wang et al., 2024), traffic management (Rehena and Janssen, 2018; Wang et al., 2021), and personalized location services (Zhao et al., 2020; Wang et al., 2020).
-Collecting trajectory data involves the coordination of multiple devices (Chen et al., 2024).
-The process begins with continuously recorded GPS coordinates from devices like smartphones or IoT sensors at regular intervals.
-These data points are then sent to central servers or databases for storage and analysis.
-However, factors such as network transmission instability, device power-saving settings, and poor coverage in high-traffic or remote areas will result in the loss of trajectory points, making the collected data sparse.
-Such sparse trajectories failed to accurately describe movement behavior and route choices, thus limiting their effectiveness in trajectory-based applications.
-To mitigate this problem, it is essential to recover the missing points from sparse trajectories which improves the integrity of trajectory information.
-Such task is usually termed trajectory recovery.
-Many works have been proposed to tackle trajectory recovery task, which can be divided into two categories: free space trajectory recovery (Chen et al., 2011, 2023a) and map-matched trajectory recovery (Wei et al., 2024; Chen et al., 2023b).
-Free space trajectory recovery aims to directly restore the coordinates of missing trajectory points but cannot ensure that the recovered trajectory is constrained to the road network.
-Typically, an additional map-matching process is required before these trajectories can be used in navigation-based applications.
-Map-matched trajectory recovery focuses on directly recovering trajectories on the road network.
-In this paper, we focus on the latter setting because it can be directly utilized by downstream tasks without the need for further processing.
-Despite the progress made by existing methods on map-matched trajectory recovery, they still face some challenges as listed below.
-1) The availability of dense trajectories is often limited.
-Deep learning-based trajectory recovery models are inherently data-driven, requiring a substantial number of dense trajectories as labels for supervised training.
-However, due to the aforementioned factors, most collected and stored trajectories are sparse, resulting in a limited availability of large-scale dense trajectory data.
-As a result, the lack of sufficient dense trajectories for training makes existing deep learning models prone to overfitting and leads to diminished performance.
-2) Existing models struggle to generalize to sparse trajectories with varying sampling intervals.
-Sparse trajectories often exhibit inconsistent sampling intervals due to network instability.
-However, most trajectory recovery models are trained on sparse trajectories with fixed sampling intervals, limiting their ability to generalize to trajectories with varying intervals.
-As a result, when confronted with trajectories that have different sampling intervals, these models require separate training for each interval to recover trajectories effectively, leading to significantly increased training costs.
-Figure 1. An example of road network evolution process.
-3) External factors related to the recovery of missing points are not incorporated.
-External factors, such as road conditions, provide valuable insights that can enhance trajectory recovery.
-For instance, in the scenario illustrated in Figure 1, given the observed points $p_{1}$ and $p_{4}$ and the road conditions at time $t_{1}$, it is more likely that the user took route $R_{1}$ rather than $R_{2}$.
-Additionally, based on the road conditions at times $t_{2}$ and $t_{3}$, as shown in Figures 1 (b) and (c), it becomes easier to infer that the user first accelerated and then decelerated between $p_{1}$ and $p_{4}$.
-However, existing methods only consider information from the known trajectory points, overlooking the significance of external factors.
-To address the aforementioned challenges, we propose a Pre-trained Language Model for Trajectory Recovery (PTR).
-By leveraging a pre-trained language model (PLM) that has been trained on large-scale corpus datasets and fine-tuning it with a limited set of dense trajectories, PTR can effectively recover trajectories even when only scarce dense trajectory data is available.
-PTR incorporates two key components to generalize to sparse trajectories with varying sampling intervals: first, an explicit trajectory prompt, which uses natural language to describe both the trajectory recovery task and its characteristics.
-This activates the PLM’s ability to recover trajectories and recognize the trajectory’s sampling interval.
-Second, a joint training strategy, where PTR is trained on trajectory data with multiple sampling intervals, further enhancing its generalization capabilities.
-To account for external factors, we introduce an implicit trajectory prompt, which integrates rich contextual information, such as road conditions, to provide insights on missing points.
-Lastly, we propose a trajectory embedder that encodes both observed and missing trajectory points, transforming them into representations interpretable by PLMs.
-This allows the PLM to capture the spatiotemporal correlations within the trajectory.
+A trajectory is a sequence of timestamped locations that describe the movement of individuals or vehicles, represented as $\mathcal{T}=\langle p_{1},\dots,p_{|\mathcal{T}|}\rangle$ of length $|\mathcal{T}|$, where $p_{i}=(lat_{i},lng_{i},t_{i})$ denotes the latitude, longitude, and timestamp of the $i$-th point.
+The sampling interval of $\mathcal{T}$ is $\mu=t_{i}-t_{i-1},\forall i\in\{2,\dots,|\mathcal{T}|\}$.
+With the advances in GPS devices and geopositioning technologies, a massive amount of trajectory data has been collected.
+These data play a pivotal role in various applications, such as urban planning (Wang et al., 2024), traffic management (Rehena and Janssen, 2018; Wang et al., 2021), and personalized location services (Zhao et al., 2020; Wang et al., 2020).
+However, factors such as network instability, device malfunctions, or cost-saving settings often lead to missing trajectory points, making the data sparse with large sampling intervals (Ren et al., 2021; Zhao et al., 2019).
+Sparse trajectories fail to accurately reflect movement behavior and route choices, limiting their usefulness.
+To address this issue, recovering the missing points in sparse trajectories is crucial for preserving trajectory completeness—a task usually referred to as trajectory recovery.
+Many works have been proposed to tackle trajectory recovery, broadly categorized into free space trajectory recovery (Chen et al., 2011, 2023a) and map-matched trajectory recovery (Wei et al., 2024; Chen et al., 2023b).
+Free space methods directly predict the coordinates of missing points but do not ensure alignment with the road network, typically requiring a subsequent map-matching step.
+Map-matched methods, by contrast, aim to recover missing points directly on the road network, making them more convenient for downstream applications.
+Despite the progress of existing methods, map-matched trajectory recovery still faces the following challenges:
+1) Limited dense trajectory data.
+Trajectory recovery models are data-driven and typically require large-scale pairs of sparse and dense trajectories.
+However, because of cost-saving settings and device malfunctions, most collected trajectories remain sparse.
+The limited availability of dense trajectory data makes existing models prone to overfitting and hampers their performance.
+2) Difficulty in generalizing to varying sampling intervals.
+Due to network instability and device malfunctions, a sparse trajectory dataset often contains trajectories with a mixture of different sampling intervals (Wang et al., 2021).
+Existing works (Wei et al., 2024; Chen et al., 2023b; Ren et al., 2021) usually treat all intervals in the same way, overlooking the varying spatiotemporal correlations of sparse trajectories entangled with multiple sampling intervals (Chen et al., 2024; Che et al., 2018).
+As a result, when encountering trajectories with unseen sampling intervals, these models often require retraining, which increases costs.
+Figure 1. Impact of road conditions on route selection and movement patterns.
+3) Non-trivial extraction of dynamic road conditions for missing points.
+Road conditions are key to understanding movement patterns.
+Existing work (Wei et al., 2024) incorporates road conditions around observed points but ignores those at missing points.
+For example, in Figure 1(b) and (c), knowing the conditions at $p_{2}$ and $p_{3}$ reveals that the user is gradually decelerating on $R_{1}$ due to congestion.
+If we only consider conditions at $p_{1}$ and $p_{4}$, we might infer that the user takes $R_{1}$ instead of $R_{2}$ (Figure 1(a)), but not the finer details of the movement.
+Since the exact locations of missing points are unknown, their road conditions cannot be directly extracted.
+To tackle these challenges, we propose Pre-trained Language Model for Trajectory Recovery (PLMTrajRec).
+We leverage a pre-trained language model (PLM) that has already learned useful representations from large-scale corpus data and fine-tune it using only a small number of dense trajectories.
+This enables PLMTrajRec to recover missing points even when dense trajectories are scarce (Challenge 1).
+To handle sparse trajectories with varying sampling intervals (Challenge 2), we introduce an interval and feature-guided (IF-guided) explicit trajectory prompt.
+It encodes both sampling intervals and movement features into the PLM through natural language, helping the model identify these intervals and capture key trajectory characteristics.
+We then propose an interval-aware trajectory embedder to standardize different sampling intervals and learn their spatiotemporal correlations.
+To infer road conditions for missing points (Challenge 3), we design an area flow-guided (AF-guided) implicit trajectory prompt that gathers traffic flows in each region.
+We also present a road condition passing mechanism that uses road conditions from nearby observed points to estimate those of the missing points.
 Overall, the main contributions of our work can be summarized as follows:
-  • We introduce PTR, a novel framework for trajectory recovery that utilizes PLMs to efficiently recover trajectories using a limited set of dense trajectories.
-  • We propose explicit trajectory prompts that enable the PLM to accurately interpret the trajectory recovery task and determine the sampling intervals of the trajectories.
-Combined with a joint training strategy on trajectories with various sampling intervals, PTR generalizes effectively to sparse trajectories across different intervals.
-  • We present an implicit trajectory prompt to model external factors influencing missing trajectory points and a trajectory embedder to transform the trajectory representation into a form that the PLM can process.
-  • We conduct extensive experiments on two real-world trajectory datasets, each with three different sampling intervals.
-The results show that our proposed model achieves state-of-the-art performance in both accuracy and efficiency.
+  • We introduce PLMTrajRec, a trajectory recovery model with strong scalability.
+By leveraging a pre-trained language model, PLMTrajRec can be fine-tuned on a small set of dense trajectories while still achieving effective recovery.
+  • We propose an IF-guided explicit trajectory prompt and an interval-aware trajectory embedder to help PLMTrajRec generalize to sparse trajectories with various sampling intervals.
+The prompt encodes intervals and movement features into the PLM, and the embedder unifies intervals and learns spatiotemporal correlations.
+  • We design an AF-guided implicit trajectory prompt to capture road conditions by aggregating traffic flows, and a road condition passing mechanism to infer missing-point conditions from nearby observations.
+  • We conduct extensive experiments on two real-world datasets, each with three sampling intervals, showing that PLMTrajRec achieves state-of-the-art performance in effectiveness, scalability, and generalizability.
 2. Related Work
-2.1. Free-space Trajectory Recovery
-Trajectory recovery in free space aims to directly restore the GPS coordinates of missing points.
-Early approaches relied on predefined rules to model mobile objects (Chen et al., 2016; Su et al., 2013).
-For example, Linear (Hoteit et al., 2014) assumes that vehicle movement follows uniform linear motion and employs linear interpolation for trajectory recovery.
-MPR (Chen et al., 2011) assumes vehicles travel along the most frequently used routes, utilizing historical trajectory data to determine route frequencies for recovery.
-Other studies (Banerjee et al., 2014; Zheng et al., 2012) use Markov models to capture spatial transitions between trajectory points.
-However, these methods are limited to modeling low-order transitions and struggle to capture the global spatial-temporal dependencies of trajectories.
-Recent deep learning-based approaches have provided more effective solutions for trajectory recovery (Xu et al., 2017, 2022).
-DHTR (Wang et al., 2019) employs a GRU-based seq2seq model to analyze user route transitions, integrating Kalman filtering to improve recovery accuracy.
-AttnMove (Xia et al., 2021) uses attention mechanisms and LSTM to capture overall human movement preferences.
-PeriodicMove (Sun et al., 2021) models each user trajectory as a graph, employing graph neural networks to efficiently capture user preferences by leveraging historical route transitions.
-TERI (Chen et al., 2023a) introduces a two-stage framework based on transformers to recover trajectories with unequal spacing, while TrajBERT (Si et al., 2023) utilizes the BERT (Devlin et al., 2018) framework to recover implicit trajectories, focusing on coarser-grained outputs such as area or Wi-Fi IDs.
-Although these methods show promising results, they require an additional map-matching process to align the recovered trajectories with the road network before they can be applied in navigation systems.
-This two-stage process leads to error accumulation and introduces significant time overhead, as the map-matching step is time-consuming.
-2.2. Map-matched Trajectory Recovery
+Free-space Trajectory Recovery.
+Trajectory recovery in free space aims to directly restore missing GPS coordinates.
+Early methods relied on predefined rules to model mobile objects (Chen et al., 2016; Su et al., 2013), such as the vehicle moving with uniform linear motion (Hoteit et al., 2014) or taking the most popular route (Chen et al., 2011).
+Other approaches (Banerjee et al., 2014; Zheng et al., 2012) use Markov models to capture spatial transitions.
+However, these methods are limited to modeling low-order transitions and fail to capture the global spatial-temporal dependencies essential for accurate trajectory recovery.
+Recent deep learning-based methods have provided more effective solutions (Xu et al., 2017, 2022).
+DHTR (Wang et al., 2019) employs a GRU-based seq2seq model to analyze user route transitions, integrating Kalman filtering to improve accuracy.
+AttnMove (Xia et al., 2021) leverages attention mechanisms and LSTM to capture movement preferences, PeriodicMove (Sun et al., 2021) models trajectories as graphs using GNNs to extract user patterns.
+TERI (Chen et al., 2023a) employs a transformer-based two-stage framework for irregular interval trajectory recovery, and TrajBERT (Si et al., 2023) adapts BERT (Devlin et al., 2018) for implicit trajectory recovery at a coarser granularity.
+Despite their effectiveness, these methods require an additional map-matching step to align recovered trajectories with road networks before they can be applied in navigation systems.
+This two-stage process leads to error accumulation and significant time overhead due to the computational cost of map-matching.
+Map-matched Trajectory Recovery.
 Compared with the above methods, map-matched trajectory recovery incorporates the road network as input and aims to recover the trajectory directly onto the road.
 To this end, MTrajRec (Ren et al., 2021) first represents the trajectory point with a road segment and moving rate, utilizing the seq2seq-based multi-task framework to capture the spatiotemporal correlations within the trajectory.
-Following this idea, RNTrajRec (Chen et al., 2023b) further models the relationship between trajectories and road networks, designing a spatial-temporal transformer architecture to encode sparse trajectories.
+Following this idea, RNTrajRec (Chen et al., 2023b) models the relationship between trajectories and road networks, designing a spatial-temporal transformer architecture to encode sparse trajectories.
+LightTR (Liu et al., 2024) proposes a lightweight trajectory recovery framework based on federated learning.
 MM-STGED (Wei et al., 2024) models sparse trajectories from a graph perspective, considering both micro and macro semantic information of trajectories.
-Although these methods demonstrate promising progress in trajectory recovery, their performance are still hindered by the poor availability of large-scale dense trajectories and limited in generalization to different sampling intervals discussed in Section 1.
+Although they demonstrate promising progress in trajectory recovery, their performance is still hindered by the poor availability of large-scale dense trajectories and limited in generalization to different sampling intervals discussed in Section 1.
 3. Preliminaries
@@
 The map-matched trajectory is represented as $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{|\mathcal{T}_{m}|}\rangle$, where each $q_{j}=(e_{j},r_{j},t_{j})$ signifies the vehicle’s location at time $t_{j},j\in\{1,\cdots,|\mathcal{T}_{m}|\}$.
-Here, $e\in\mathcal{V}$ is the matched road segment, and $r$ is the moving ratio, which quantifies the proportion of the distance traveled from the starting point of a road segment in relation to the total length of that road segment.
+Here, $e\in\mathcal{E}$ is the matched road segment, and $r$ is the moving ratio, which quantifies the proportion of the distance traveled from the starting point of a road segment in relation to the total length of that road segment.
 Figure 2. An illustration of map-matched trajectory, road segment $e$, and moving ratio $r$.
@@
 Therefore, the moving ratio $r_{6}$ can be calculated as $r_{6}=\frac{d_{6}}{l_{3}}$.
-Figure 3. The framework of PTR, consists of Trajectory Prompt, Trajectory Embedder, and PLM-based Trajectory Encoder.
+Figure 3. The framework of PLMTrajRec, consists of Dual Trajectory Prompts, Interval-aware Trajectory Embedder, and PLM-based Trajectory Encoder.
 3.2. Problem Statement
@@
 4.1. Overall Pipeline
-In this paper, we present PTR, which fine-tunes a PLM pre-trained on a large-scale corpus to efficiently recover trajectories with limited dense data.
-The framework of PTR, shown in Figure 3, comprises three main components: trajectory prompts, a trajectory embedder, and a PLM-based trajectory encoder.
-Trajectory prompts offer essential information from two key components.
-First, explicit trajectory prompts, expressed in natural language, outline the trajectory recovery task and key trajectory characteristics.
-These prompts activate the PLM’s ability to recover trajectories and identify varying sampling intervals in sparse trajectories.
-With joint training on trajectories across different intervals, PTR achieves robust generalization.
-Second, implicit trajectory prompts model road conditions, providing crucial external factors of missing points.
-The trajectory embedder effectively enhances the PLM’s understanding of trajectories.
-It begins by normalizing sparse trajectories with varying sampling intervals $\mu$ into a uniform interval $\epsilon$.
-Next, each trajectory point, including both observed and missing points, is encoded into an embedding that captures both local and global semantics.
-Finally, the embedding is transformed into a format understood with the PLM.
-In the PLM-based trajectory encoder, most of the encoder’s parameters are frozen to preserve the pre-trained knowledge, while the multi-head attention layer remains trainable, allowing the model to capture the spatiotemporal correlations within the trajectory.
+In this paper, we present both scalable and generalizable trajectory recovery model, PLMTrajRec, by fine-tuning a PLM that is pre-trained on a large-scale corpus with limited dense trajectory data.
+The framework of PLMTrajRec, shown in Figure 3, comprises three main components: dual trajectory prompts, an interval-aware trajectory embedder, and a PLM-based trajectory encoder.
+The dual trajectory prompts provide essential information through two key components.
+First, the interval and feature-guided (IF-guided) explicit trajectory prompt, expressed in natural language, injects the sampling interval of sparse trajectories and their movement features into PLM.
+This enhances the model’s ability to recognize the sampling intervals of each sparse trajectory and capture trajectory characteristics.
+Second, area flow-guided (AF-guided) implicit trajectory prompts model road conditions, offering crucial contextual information to improve the recovery of missing trajectory points.
+The interval-aware trajectory embedder first normalizes sparse trajectories with varying sampling intervals $\mu$ into a unified interval $\epsilon$, effectively handling the diverse spatiotemporal correlations associated with different sampling intervals and enhancing model generalization.
+Subsequently, each trajectory point, including both observed and missing points, is encoded into an embedding that the PLM can process.
+In the PLM-based trajectory encoder, pre-trained BERT is employed to capture information bi-directionly, and most of the parameters are frozen to preserve the pre-trained knowledge, while the multi-head attention layer remains trainable, allowing the model to capture the spatiotemporal correlations within the trajectory.
 Finally, the output layer predicts the road segment $e$ and the moving ratio $r$ for the recovered map-matched trajectory at each step.
-4.2. Trajectory Prompts
-To enable PLMs to recover trajectories with different intervals and efficiently model external factors for missing trajectory points, we introduce explicit and implicit trajectory prompts.
-4.2.1. Explicit Trajectory Prompt
-The explicit trajectory prompt offers a structured textual description at both the task and trajectory levels, enabling PTR to adapt to varying sampling intervals.
-At the task level, it includes the following components:
+4.2. Dual Trajectory Prompts
+To enable PLM to recover trajectories with different intervals and effectively model road conditions for missing trajectory points, we introduce dual trajectory prompts, including an interval and feature-guided (IF-guided) explicit trajectory prompt and an area flow-guided (AF-guided) implicit trajectory prompt.
+4.2.1. IF-guided Explicit Trajectory Prompt
+The IF-guided explicit trajectory prompt offers a structured textual description of the sampling intervals and movement features of sparse trajectories, enabling PLMTrajRec to identify varying sampling intervals and capture trajectory features.
+The prompts related to the sampling interval include the following components:
   • ¡Task Part¿: Sparse trajectory recovery.
@@
 […]
-At the trajectory level, the explicit trajectory prompt provides detailed characteristics of the trajectory, consisting of:
+The prompts related to the movement features include the following contents:
 […]
@@
 The ¡Movement Part¿ supports the PLM in inferring the trajectory’s movement.
+We give a detailed example of the IF-guided explicit trajectory prompt in Appendix C.
 After obtaining the prompt for each part, we use the tokenizer and word token embedding in PLMs to convert text into embeddings.
 The embedding of ¡Task Part¿, ¡Target Part¿, ¡Content Part¿, ¡Time Part¿, and ¡Movement Part¿ are denoted by $\textbf{H}_{ta}$, $\textbf{H}_{tar}$, $\textbf{H}_{c}$, $\textbf{H}_{t}$, and $\textbf{H}_{m}$.
-These embeddings are then concatenated to form the overall explicit trajectory prompt embedding.
+These embeddings are then concatenated to form the overall IF-guided explicit trajectory prompt embedding.
 (1) $\textbf{H}^{e}=\textbf{H}_{ta}||\textbf{H}_{tar}||\textbf{H}_{c}||\textbf{H}_{t}||\textbf{H}_{m},$
 where ||||| | is the concatenate operation.
-4.2.2. Implicit Trajectory Prompt
-We incorporate road conditions as external factors for missing trajectory points, as they provide insights into both the surrounding environment and the object’s movement, which are all helpful for trajectory recovery.
+4.2.2. AF-guided Implicit Trajectory Prompt
+Road conditions can provide insights into both the surrounding environment and the object’s movement, which are helpful for trajectory recovery.
 For instance, vehicles typically slow down in congested areas and accelerate in less congested regions.
-Given the complexity and variability of real-world road conditions, which are difficult to express in natural language, we represent them as an implicit trajectory prompt.
+Given the complexity and variability of real-world road conditions are difficult to express in natural language, we represent them as the implicit trajectory prompt.
 To obtain the road conditions for each trajectory point, we first calculate the average road conditions across all areas and time intervals, then extract the relevant information for each point.
@@
 For each grid cell, we compute the total traffic, resulting in a regional flow matrix $\textbf{RC}\in\mathbb{R}^{I\times J\times T}$.
-To capture the spatiotemporal correlations among different regions, we apply a 2D convolution (2D CNN) along the spatial dimensions and a 1D CNN along the temporal dimension, yielding the road condition representation $\textbf{H}_{rc}\in\mathbb{R}^{I\times J\times T\times F}$, where $F$ is the number of features.
+Each element in RC represents the number of traffic flows at region $i,j$ at time $t$.
+To capture the spatiotemporal correlations among different regions, we apply a 2D convolution (2D CNN) along the spatial dimension and a 1D CNN along the temporal dimension, yielding the road condition representation $\textbf{H}_{rc}\in\mathbb{R}^{I\times J\times T\times F}$, where $F$ is the number of features.
 Formally:
 (2) $\mathbf{H}_{rc}=\text{Conv1d(Conv2d({RC}))}\in\mathbb{R}^{I\times J\times T\times F}$
-Obviously, the road conditions of a trajectory point are indexed by its longitude, latitude, and timestamp.
+Obviously, extracting the road conditions of a trajectory point needs its longitude, latitude, and timestamp.
 However, it is challenging for missing points as their geographic coordinates are unknown.
-To address this, we introduce a road condition diffusion strategy that approximates the missing conditions using neighboring traffic.
+Inspired by the message passing mechanism (Gilmer et al., 2017), we introduce a road condition passing mechanism that approximates the road conditions of missing points by leveraging the surrounding road conditions.
 The implementation details will be elaborated in Section 4.3.2.
 Figure 4. An illustration of trajectory preprocessing layer.
-4.3. Trajectory Embedder
-To efficiently encode trajectory points and enable PLMs to better understand them, we introduce the trajectory embedder.
-4.3.1. Trajectory Preprocessing Layer
-For different sparse trajectories, their sampling intervals $\mu$ are not constant, such as 4 minutes, 2 minutes, or 1 minute.
-To convert these into a uniform sampling interval, we introduce a placeholder ‘[m]’ to mark missing trajectory points and generate a preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, where the length of $|\mathcal{T}^{\epsilon}|=\frac{p_{|\mathcal{T}_{s}|}.t-p_{1}.t}{\epsilon}$.
+4.3. Interval-aware Trajectory Embedder
+To make the model deal with sparse trajectories with different sampling intervals and capture the spatiotemporal correlation, we propose an interval-aware trajectory embedder.
+4.3.1. Trajectory Interval Unification
+For different sparse trajectories, their sampling intervals $\mu$ are not constant, such as 4 minutes, 2 minutes, or 1 minute, which bring different spatiotemporal correlations between trajectory points.
+To this end, we convert these into the same sampling interval with the target trajectory by introducing a placeholder ‘[m]’ to mark missing trajectory points and generate a preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, where the length of $\mathcal{T}^{\epsilon}$ is $\frac{p_{|\mathcal{T}_{s}|}.t-p_{1}.t}{\epsilon}+1$.
 Although the location of ‘[m]’ is unknown, its timestamp can still be calculated, as illustrated in Figure 4.
 4.3.2. Trajectory Feature Extractor
-To embed trajectory points, we propose the trajectory feature extractor.
-Given the preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, there are two cases for its trajectory point characteristics:
+Given the preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, there are two cases for extract trajectory point features:
   • Case 1: the trajectory point $s\in\mathcal{T}^{\epsilon}$ is observed. i.e., $\exists\,k\in\{1,\cdots,|\mathcal{T}_{s}|\},p_{k}.t=s.t$.
   • Case 2: the location of trajectory point $s\in\mathcal{T}^{\epsilon}$ is missing. i.e., $s=[m]$, note that its time $s.t$ is known.
-For case 1, we use Learnable Fourier Features (Tancik et al., 2020; Li et al., 2021) to encode the continuous latitude and longitude of $s$ into an $F$-dimensional vector using the feature mapping function $\Phi(x):\mathbb{R}\to\mathbb{R}^{F}$.
+For case 1, we use Learnable Fourier Features (LFF) (Tancik et al., 2020; Li et al., 2021) to encode the continuous latitude and longitude of $s$ into an $F$-dimensional vector using the feature mapping function $\Phi(x):\mathbb{R}\to\mathbb{R}^{F}$.
 Formally:
@@
 where $W_{r}\in\mathbb{R}^{F/2}$ and $W_{\Phi}\in\mathbb{R}^{F\times F}$ represent the learnable parameters, and $x\in\{s.lat,s.lon\}$.
-Using the feature mapping function $\Phi(\cdot)$, the relative information between points $x$ and $y$ can be captured through multiplication operations.
-This relative information helps infer crucial details like the distance between trajectory points, which is important for understanding movement.
-For instance, a larger distance between two points may indicate a higher likelihood of vehicle acceleration.
+Using the feature mapping function $\Phi(\cdot)$, the relative information $x-y$ between points $x$ and $y$ can be captured through multiplication operations, which is important for understanding movement, such as distance.
+We provide detailed insight and proof of LFF in Appendix D.
 Additionally, considering that vehicles move within the road network, the relationship between a trajectory point and its surrounding road segments is critical.
@@
 Then, we obtain the road network representation $\mathbf{h}_{s}^{\text{road}}$ by:
-(5) $\mathbf{h}_{s}^{\text{road}}=\frac{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})*\textbf{M}_{l}}{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})},$
+(5) $\mathbf{h}_{s}^{\text{road}}=\frac{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})\cdot\textbf{M}_{l}}{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})},$
 where $\textbf{M}_{l}\in\mathbb{R}^{F}$ is the learnable embedding of road segment $l$, and $\mathcal{|V|}$ is the total number of road segments.
@@
 (6) $\mathbf{h}_{s}=W_{1}[(\Phi(s.lat)+\Phi(s.lng))||\mathbf{h}_{s}^{\text{road}}]+b_{1},$
-where $W_{1}\in\mathbb{R}^{2F\times F}$ and $b_{1}\in\mathbb{R}^{F}$ is the learnable parameters.
+where $W_{1}\in\mathbb{R}^{F\times 2F}$ and $b_{1}\in\mathbb{R}^{F}$ is the learnable parameters.
 For case 2, where the location of point $s$ is unknown, here we use road conditions to represent its features, as described in Section 4.2.2.
-Since the road conditions at a specific point are influenced by the surrounding conditions, which propagate along both temporal and spatial dimensions, we propose a road condition diffusion strategy to model the road conditions of the missing location.
+Since the road conditions at a specific point are influenced by the surrounding conditions, which propagate along both temporal and spatial dimensions, we propose a road condition passing mechanism to infer the road conditions of the missing location inspired by the message passing mechanism (Gilmer et al., 2017).
+Road Condition Passing Mechanism.
 First, for the missing point $s$, we identify the observed forward and backward trajectory points, $s_{f}$ and $s_{b}$, respectively.
@@
 (7) $\begin{split}\mathbf{h}_{rc}^{f}&=\mathbf{H}_{rc}[\pi_{lat}(s_{f}.lat),\pi_{lng}(s_{f}.lng),\pi_{t}(s_{f}.t)],\\ \mathbf{h}_{rc}^{b}&=\mathbf{H}_{rc}[\pi_{lat}(s_{b}.lat),\pi_{lng}(s_{b}.lng),\pi_{t}(s_{b}.t)],\end{split}$
-where $\pi_{lat}$, $\pi_{lng}$, and $\pi_{t}$ are index functions mapping latitude, longitude, and time to their respective indices.
+where $\pi_{lat}$, $\pi_{lng}$, and $\pi_{t}$ are index functions mapping latitude, longitude, and time to their corresponding indices.
 Next, we calculate the time interval between $s$ and $s_{f}$ and $s_{b}$, denoted as $\triangle t_{f}=s.t-s_{f}.t$ and $\triangle t_{b}=s_{b}.t-s.t$.
@@
 In addition, we encode the time intervals $\triangle t_{f}$ and $\triangle t_{b}$ to quantify the relative position of $s$. The final representation of $s$ is given by:
-(9) $\begin{split}\mathbf{h}_{s}&=W_{2}[\textbf{m}\text{ }||\text{ FC}(\triangle t_{f}||\triangle t_{b})\text{ }||\text{ }\mathbf{h}_{rc}^{s}]+b_{2},\\ \text{FC}(\cdot)&:\mathbb{R}^{2}\to\mathbb{R}^{F},\end{split}$
-where $\textbf{m}\in\mathbb{R}^{F}$ is a learnable vector to represent the location is missing, and FC()($\cdot$)( ⋅ ) is the fully connection layer.
-$W_{2}\in\mathbb{R}^{3F\times F}$ and $b_{2}\in\mathbb{R}^{F}$ are the learnable parameters.
+(9) $\mathbf{h}_{s}=W_{2}[\textbf{m}\text{ }||\text{ FC}(\triangle t_{f}||\triangle t_{b})\text{ }||\text{ }\mathbf{h}_{rc}^{s}]+b_{2},$
+where $\textbf{m}\in\mathbb{R}^{F}$ is a learnable vector to represent the location is missing, and $\text{FC}(\cdot):\mathbb{R}^{2}\to\mathbb{R}^{F}$ is the fully connection layer.
+$W_{2}\in\mathbb{R}^{F\times 3F}$ and $b_{2}\in\mathbb{R}^{F}$ are the learnable parameters.
 Finally, the overall trajectory representation is $\textbf{H}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times F}$.
@@
 4.4. PLM-based Trajectory Encoder
-As shown in Figure 3, we use the stacked Transformer layers as the fundamental architecture of PLMs.
-To make PLMs fully adaptable to the trajectory recovery task, we employ the Low-Rank Adaptation (LoRA) algorithm (Hu et al., 2021).
+We use the pre-trained BERT as the fundamental architecture of PLMs, considering that its encoder-only structure is well-suited to the reconstruction task due to its effective use of bi-directional contextual information from the trajectory (Sun et al., 2023).
+To make PLM fully adaptable to the trajectory recovery task, we employ the Low-Rank Adaptation (LoRA) algorithm (Hu et al., 2021).
 LoRA introduces additional parameters within the Transformer block, reducing the number of trainable parameters and computational complexity while enhancing the PLM’s performance on new tasks.
@@
 (13) $\textbf{Z}^{l}=\text{LoRA(}\text{Transformer }(\textbf{Z}^{l-1}))$
-After stacking $L$ Transformer layers, we discard the explicit trajectory prompt portion and obtain the output embeddings $\mathbf{Z}^{\prime}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times d}$.
+After stacking $L$ Transformer layers, we discard the IF-guided explicit trajectory prompt portion and obtain the output embeddings $\mathbf{Z}^{\prime}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times d}$.
 For the $j$-th element of $\textbf{Z}^{\prime}$, we apply the softmax function to calculate the probability of each road segment and use the argmax function to return the predicted road segment $e_{j}$.
@@
 4.5.1. Loss Function
-We use multi-task learning to simultaneously optimize road segment recovery and moving ratio recovery tasks. For the road segment recovery, we use the cross-entropy as the loss function:
+We use multi-task learning to optimize both road segment recovery and moving ratio recovery simultaneously. For the road segment, we use the cross-entropy loss function:
 (14) $\mathcal{L}_{e}=-\frac{1}{|\mathcal{T}^{\epsilon}|}\sum_{j=1}^{|\mathcal{T}^{\epsilon}|}\mathrm{log}(\hat{e}_{j}|\textbf{Z}^{\prime}_{j})$
-For the moving ratio, we use the mean squared error as the loss function:
+For the moving ratio, we use the mean squared error loss function:
 (15) $\mathcal{L}_{r}=\frac{1}{|\mathcal{T}^{\epsilon}|}\sum_{j=1}^{|\mathcal{T}^{\epsilon}|}|r_{j}-\hat{r}_{j}|^{2}$
 The final loss function is $\mathcal{L}=\mathcal{L}_{e}+\lambda\mathcal{L}_{r}$, where $\lambda$ is a hyper-parameter to balance two tasks.
-4.5.2. Training Strategy
-To generalize PTR in recovering sparse trajectories with various sampling intervals, we propose the following training approach.
+4.5.2. Joint Training Strategy
+To generalize PLMTrajRec in recovering sparse trajectories with various sampling intervals, we propose the following training approach.
 For the dense trajectory dataset $\mathbb{T}$, which is set as a 15-second sampling interval in our experiments.
-To create sparse trajectory datasets, we re-sample $\mathbb{T}$ at three different intervals: 1 minute, 2 minutes, and 4 minutes, resulting in three datasets: $\mathbb{T}_{1}$, $\mathbb{T}_{2}$, and $\mathbb{T}_{4}$.
-These are then combined to form a larger training dataset, $\mathbb{T}_{all}$.
+To create sparse trajectory datasets, we re-sample each trajectory $\mathcal{T}\in\mathbb{T}$ at three different intervals: 1 minute, 2 minutes, and 4 minutes, resulting in three datasets: $\mathbb{T}_{1}$, $\mathbb{T}_{2}$, and $\mathbb{T}_{4}$.
+These are then combined to form a larger training dataset $\mathbb{T}_{all}$.
+Formally,
 (16) $\mathbb{T}_{all}=\mathbb{T}_{1}\cup\mathbb{T}_{2}\cup\mathbb{T}_{4}.$
-By training PTR on $\mathbb{T}_{all}$, we achieve balanced performance across the different sampling intervals.
+By training PLMTrajRec on $\mathbb{T}_{all}$, we achieve balanced performance across the different sampling intervals.
 Subsequently, we fine-tune the model on each individual sparse trajectory dataset to further improve trajectory recovery performance.
 5. Experiments
-In this section, we present comprehensive experiments to evaluate the effectiveness of PTR on two real-world trajectory datasets with three different sparse trajectory sampling intervals.
-These experiments are designed to answer the following key questions:
-  • RQ1: How does PTR’s performance compare to other baseline methods across different sampling intervals of sparse trajectories?
-  • RQ2: How well does PTR perform when trained with limited data?
-  • RQ3: What is the contribution of each component in PTR to the overall model performance?
+In this section, we present comprehensive experiments to evaluate the effectiveness, scalability, and generalization capabilities of PLMTrajRec on two real-world trajectory datasets with three different sparse trajectory sampling intervals.
 5.1. Datasets
@@
 The detailed overview of the dataset characteristics is provided in Table 1.
+Table 1. Dataset Description.
+Types Chengdu Porto
+Sampling interval 15s 15s
+# Trajectory 118,354 322,079
+# Road segment 2504 2224
+Latitude range 30.655 $\sim$ 30.727 41.142 $\sim$ 41.174
+Longitude range 104.043 $\sim$ 104.129 -8.652 $\sim$ -8.578
+Length / Width 8.22 km / 8.00 km 6.19 km / 3.56 km
 5.2. Baselines
-To evaluate the effectiveness of our model, we compare PTR with ten baseline methods.
-These include five free space trajectory recovery models: HMM (Newson and Krumm, 2009) + ShortestPath, Linear (Hoteit et al., 2014) + HMM (Newson and Krumm, 2009), MPR (Chen et al., 2011) + HMM (Newson and Krumm, 2009), DHTR (Wang et al., 2019) + HMM (Newson and Krumm, 2009), and AttnMove (Xia et al., 2021) + Rule, as well as five map-matched trajectory recovery models: MTrajRec (Ren et al., 2021), T2vec (Li et al., 2018) + Decoder, T3s (Yang et al., 2021) + Decoder, RNTrajRec (Chen et al., 2023b), and MM-STGED (Wei et al., 2024).
-The details of each baseline are provided in Appendix B.
+To evaluate the effectiveness of our model, we compare PLMTrajRec with 12 baseline methods.
+These include five free space trajectory recovery models: HMM (Newson and Krumm, 2009) + ShortestPath, Linear (Hoteit et al., 2014) + HMM (Newson and Krumm, 2009), MPR (Chen et al., 2011) + HMM (Newson and Krumm, 2009), DHTR (Wang et al., 2019) + HMM (Newson and Krumm, 2009), and AttnMove (Xia et al., 2021) + Rule, as well as seven map-matched trajectory recovery models: MTrajRec (Ren et al., 2021), T2vec (Li et al., 2018) + Decoder, T3s (Yang et al., 2021) + Decoder, TERI (Chen et al., 2023a) + Decoder, TrajBERT (Si et al., 2023) + Decoder, RNTrajRec (Chen et al., 2023b), and MM-STGED (Wei et al., 2024).
+Due to space constraints, the details of each baseline are provided in Appendix A.
 5.3. Evaluation Metrics
@@
 For road segment recovery, we utilize Accuracy (Acc), Recall, and Precision (Prec) metrics to evaluate.
-A higher result indicates a more precise road segment recovery.
 And employ Mean Absolute Error (MAE) and the Root Mean Square Error (RMSE) in assessing the recovered GPS coordinates.
-Lower values of these metrics indicate a higher level of accuracy in the recovery results.
-The details of each metric are provided in Appendix C.
+The details of each metric are provided in Appendix B.
 Table 2. Performance comparison on two datasets with sampling intervals at 4 minutes, 2 minutes, and 1 minute, respectively.
@@
 T2vec + Decoder 66.51 75.68 78.27 307.5 915.2 53.13 60.27 77.62 256.3 571.0
+TERI + Decoder 66.42 75.59 78.36 309.2 903.8 53.59 60.02 78.26 253.9 558.3
+TrajBERT + Decoder 66.09 75.38 78.59 310.7 911.4 52.98 60.12 77.93 251.8 560.1
 RNTrajRec 67.66 75.59 79.97 306.1 886.0 54.59 60.42 79.20 248.1 549.1
 MM-STGED 70.64 76.04 81.63 266.2 829.7 57.30 59.48 80.21 222.8 510.4
-PTR 74.12 79.63 86.46 262.8 483.0 57.61 59.15 82.19 200.9 376.9
-PTR + FT 74.58 80.09 86.63 253.2 465.7 57.67 59.08 82.05 200.8 370.2
+PLMTrajRec 74.12 79.63 86.46 262.8 483.0 57.61 59.15 82.19 200.9 376.9
+PLMTrajRec + FT 74.58 80.09 86.63 253.2 465.7 57.67 59.08 82.05 200.8 370.2
 $\mu=2$ minutes $\to\epsilon=15$ seconds HMM + ShortestPath 33.85 47.89 48.31 754.1 826.2 27.30 40.05 46.00 647.3 747.0
@@
 T2vec + Decoder 75.69 78.86 81.68 231.6 783.6 62.24 65.77 78.97 173.9 438.0
+TERI + Decoder 75.32 78.74 81.38 239.0 823.7 62.14 65.39 79.18 178.3 443.9
+TrajBERT + Decoder 75.20 78.69 81.53 235.0 813.7 62.34 65.66 79.02 177.9 441.2
 RNTrajRec 75.80 79.35 81.86 218.5 757.0 63.39 65.84 79.25 171.3 433.9
 MM-STGED 78.14 80.06 83.58 197.2 696.0 65.69 66.15 80.74 152.5 400.8
-PTR 81.76 84.31 88.38 187.4 366.2 66.40 66.52 82.14 141.9 294.6
-PTR + FT 82.29 84.59 88.72 181.2 349.0 66.72 66.87 82.38 141.7 293.5
+PLMTrajRec 81.76 84.31 88.38 187.4 366.2 66.40 66.52 82.14 141.9 294.6
+PLMTrajRec + FT 82.29 84.59 88.72 181.2 349.0 66.72 66.87 82.38 141.7 293.5
 $\mu=1$ minute $\to\epsilon=15$ seconds HMM + ShortestPath 35.92 67.92 60.16 529.7 638.2 34.75 48.46 48.67 527.2 659.3
@@
 T2vec + Decoder 81.69 81.90 83.88 185.6 714.1 71.86 71.10 80.48 114.2 334.7
+TERI + Decoder 81.25 81.89 83.92 186.9 710.4 71.53 71.38 80.29 113.2 325.9
+TrajBERT + Decoder 81.38 81.72 83.65 183.2 710.7 71.63 71.47 80.39 112.7 329.5
 RNTrajRec 81.88 82.09 84.84 177.9 702.5 72.31 71.88 80.57 110.3 325.7
 MM-STGED 84.26 84.15 85.92 154.0 633.5 73.16 72.27 80.81 108.2 321.9
-PTR 87.17 87.99 90.52 141.4 290.8 74.42 72.78 82.82 95.6 211.7
-PTR + FT 88.15 88.82 91.04 138.6 289.1 75.35 73.87 83.28 95.2 220.3
+PLMTrajRec 87.17 87.99 90.52 141.4 290.8 74.42 72.78 82.82 95.6 211.7
+PLMTrajRec + FT 88.15 88.82 91.04 138.6 289.1 75.35 73.87 83.28 95.2 220.3
 5.4. Settings
@@
 To generate sparse trajectories, for each dense trajectory, we retain the first and last points and create three sparse versions by setting the sampling intervals $\mu=4$ minutes, 2 minutes, and 1 minute, respectively.
-During training, we integrate sparse trajectories with three different sampling intervals to train PTR.
+During training, we integrate sparse trajectories with three different sampling intervals to train PLMTrajRec.
 After training, we evaluate its performance in recovering trajectories across various sampling intervals to assess its generalizability.
 To further enhance performance, we fine-tune the trained model for each specific sampling interval, allowing for a more precise recovery.
-We employ PyTorch (Paszke et al., 2019) framework to implement PTR, with a learning rate of 1e-4 and a batch size of 64.
-BERT (Devlin et al., 2018) is selected as the foundation model for PLMs, as its encoder-only structure is well-suited to the reconstruction task due to its effective use of bi-directional contextual information from the trajectory (Ren et al., 2024).
+We employ PyTorch (Paszke et al., 2019) framework to implement PLMTrajRec, with a learning rate of 1e-4 and a batch size of 64.
+BERT-small  [^4] https://github.com/google-research/bert is selected as the foundation model for PLM with 4 transformer layers and the number of hidden state is 512.
 For road condition extraction, the area of interest is divided into a $64\times 64$ grid, and the time dimension is partitioned into hourly intervals.
@@
 All experiments are conducted on NVIDIA RTX A4000 GPUs.
-Table 2. Scalibility analysis on Chengdu dataset.
-Sampling Interval Rate 20% 40% 60% 80% 100%
+Table 3. Scalability analysis. The performance comparison on the Chengdu dataset when trained with different data ratios.
+Sampling Interval Data Ratio 20% 40% 60% 80% 100%
 […]
-$\mu=4$ minutes $\to\epsilon=15$ seconds MTrajRec 60.73 1048.0 63.58 968.5 64.73 929.4 65.39 914.6 65.79 904.4
-T3s + Decoder 60.49 1098.9 63.32 968.7 64.23 961.9 65.04 942.3 65.60 926.3
-T2vec + Decoder 62.17 966.1 64.92 946.1 64.95 972.3 65.31 946.8 66.51 915.2
-RNTrajRec 60.80 998.0 62.85 931.3 64.19 909.1 65.46 835.0 67.66 886.0
-MM-STGED 64.61 935.0 67.76 881.8 68.53 853.7 69.95 825.3 70.64 829.7
-PTR 68.30 585.3 71.57 515.6 72.74 512.5 73.56 488.3 74.12 483.0
 $\mu=2$ minutes $\to\epsilon=15$ seconds MTrajRec 69.85 919.8 72.58 907.4 73.97 902.3 74.32 891.3 74.52 885.7
@@
 MM-STGED 71.65 865.0 75.32 773.0 75.49 752.1 76.94 747.7 78.14 696.0
-PTR 76.29 452.8 79.37 411.5 80.37 395.4 81.18 377.4 81.76 366.2
-$\mu=1$ minute $\to\epsilon=15$ seconds MTrajRec 75.39 937.5 78.53 835.1 80.01 794.3 80.93 725.7 81.12 718.4
-T3s + Decoder 76.49 917.1 79.08 824.9 80.65 767.8 80.82 716.5 80.90 713.0
-T2vec + Decoder 75.89 845.4 79.09 746.8 79.41 752.3 81.21 742.2 81.69 714.1
-RNTrajRec 75.65 846.2 79.48 784.8 79.61 769.3 81.74 742.5 81.88 702.5
-MM-STGED 76.02 857.7 79.86 734.7 82.25 676.4 83.44 663.0 84.26 633.5
-PTR 81.62 397.5 84.59 352.0 85.75 328.7 86.73 305.4 87.17 290.8
-5.5. Experimental Results (RQ1)
+PLMTrajRec 76.29 452.8 79.37 411.5 80.37 395.4 81.18 377.4 81.76 366.2
+5.5. Experimental Results
 Table 2 reports a comparison of the results between our model and the baselines across various sampling intervals of sparse trajectories on both the Chengdu and Porto datasets.
@@
 Notably, it shows substantial gains in RMSE metrics, with an average reduction of 351.8 meters in Chengdu and 119.2 meters in Porto.
-This indicates PTR generalizes effectively, enabling it to accurately recover trajectories with varying sampling intervals.
-By fine-tuning at the specific sampling interval, PTR + FT, our model further improves Acc by 1.72% and 0.32%.
+This indicates PLMTrajRec generalizes effectively, enabling it to accurately recover trajectories with varying sampling intervals.
+By fine-tuning at the specific sampling interval, PLMTrajRec + FT, our model further improves Acc by 1.72% and 0.32%.
 These outstanding results can be attributed to the strong generalization capability of the PLM in processing sparse trajectories with varying sampling intervals.
-Additionally, the trajectory embedder plays a pivotal role in connecting trajectory data with the PLM, efficiently capturing spatial-temporal correlations and converting trajectory embeddings into a format that the PLM can effectively process and understand.
-We also analyze the convergence rate of PTR across different sampling intervals, as shown in Figure 5 and  6.
-It demonstrates rapid convergence within just a few epochs on both the Chengdu and Porto datasets.
-While fine-tuning enhances performance further, the model stabilizes by the 3th epoch, indicating that PTR can deliver satisfactory results within a short training period.
-5.6. Scalibility Analysis (RQ2)
-To evaluate the effectiveness of PTR in trajectory recovery when dense trajectory data is limited, we conduct scalability experiments on the Chengdu dataset.
+Additionally, the interval-aware trajectory embedder plays a pivotal role to capture spatial-temporal correlations and converting trajectory embeddings into a format that the PLM can effectively process and understand.
+5.6. Scalability Analysis
+To evaluate the effectiveness of PLMTrajRec in trajectory recovery when dense trajectory data is limited, we conduct scalability experiments on the Chengdu dataset.
 Specifically, we train the model using subsets of the training set at 20%, 40%, 60%, 80%, and 100%, and evaluate performance on the test set.
-The results are presented in Table 2.
+The results are presented in Table 3 and 7 of Appendix F.2.
 As the size of the training set increases, we observe performance improvements across all baseline models.
-Notably, with only 20% of the training data, PTR already surpasses most baselines, trailing only MM-STGED.
-With 40% of the training data, PTR outperforms all other models, demonstrating its strong generalization ability.
-This advantage underscores the practicality of PTR in real-world scenarios, where obtaining large-scale dense trajectory data is often challenging.
-Table 3. Ablation study on two datasets with sampling intervals at 2 minutes.
+Notably, with only 20% of the training data, PLMTrajRec already surpasses most baselines, trailing only MM-STGED.
+With 40% of the training data, PLMTrajRec outperforms all other models, demonstrating its strong scalability.
+This advantage underscores the practicality of PLMTrajRec in real-world scenarios, where obtaining large-scale dense trajectory data is often challenging.
+Table 4. Zero-shot study on Chengdu dataset with sampling intervals at 2 minutes.
 […]
-Chengdu w/o explicit trajectory prompt 81.04 84.08 88.07 205.8 394.2
-w/o implicit trajectory prompt 81.43 84.28 88.14 191.4 371.1
-w/o trajectory prompt 80.70 83.61 87.73 225.9 407.4
+MTrajRec 56.39 ($\downarrow$ 18.13%) 492.7 ($\downarrow$ 48.37%) 1046.2 ($\downarrow$ 15.34%)
+T3s + Decoder 63.81 ($\downarrow$ 10.81%) 353.0 ($\downarrow$ 31.39%) 980.6 ($\downarrow$ 9.68%)
+T2vec + Decoder 63.32 ($\downarrow$ 12.37%) 376.6 ($\downarrow$ 38.50%) 928.1 ($\downarrow$ 15.57%)
+RNTrajRec 64.93 ($\downarrow$ 10.87%) 339.6 ($\downarrow$ 35.70%) 912.8 ($\downarrow$ 17.07%)
+MM-STGED 68.27 ($\downarrow$ 9.90%) 291.8 ($\downarrow$ 32.42%) 862.6 ($\downarrow$ 19.31%)
+PLMTrajRec 79.62 ($\downarrow$ 2.14%) 195.3 ($\downarrow$ 4.05%) 379.6 ($\downarrow$ 3.53 %)
+5.7. Zero-shot Study on Sampling Interval
+We further analyze the generalizability of our model when the target sampling interval is absent during training.
+Specifically, PLMTrajRec is trained with a mixture of sampling intervals of 1 minute and 4 minutes and tested on a 2-minute sampling interval.
+The experimental results on the Chengdu dataset are shown in Table 4, where $\downarrow$ indicates the percentage decline compared to training with a 2-minute sampling interval.
+We observe that the accuracy of all models decreased when the target sampling interval was not included in the training data.
+However, our model shows a notably smaller performance drop.
+This robustness is attributed to the trajectory prompts and joint training strategy used by PLMTrajRec, which effectively identify the sampling interval of the target trajectories, demonstrating strong generalizability.
+In contrast, baseline models rely solely on trajectory point information to capture correlations, which struggles with generalization due to the inconsistency of temporal dynamics between the training and testing.
+5.8. Ablation Study
+To evaluate the effectiveness of the proposed components, we create six variants of PLMTrajRec.
+w/o IF-guided explicit trajectory prompt, w/o AF-guided implicit trajectory prompt, w/o dual trajectory prompt, w/o road network, w/o reference tokens, and PLMTrajRec - Randomly initialized BERT.
+The detailed description of each variant is shown in Appendix E.
+Table 5. Ablation study on Chengdu dataset with sampling intervals at 2 minutes.
+[…]
+w/o IF-guided explicit trajectory prompt 81.04 84.08 88.07 205.8 394.2
+w/o AF-guided implicit trajectory prompt 81.43 84.28 88.14 191.4 371.1
+w/o dual trajectory prompts 80.70 83.61 87.73 225.9 407.4
 w/o road network 81.28 84.22 88.04 194.5 375.6
 w/o reference tokens 80.83 83.36 86.68 236.6 437.2
-PTR 81.76 84.31 88.38 187.4 366.2
-Porto w/o explicit trajectory prompt 66.13 66.19 81.98 157.9 314.6
-w/o implicit trajectory prompt 66.18 66.22 82.01 149.6 298.1
-w/o trajectory prompt 65.52 65.89 81.51 171.0 339.4
-w/o road network 66.11 66.27 81.90 150.7 302.8
-w/o reference tokens 65.89 65.07 80.57 165.4 327.7
-PTR 66.40 66.52 82.14 141.9 294.6
-5.7. Ablation Study (RQ3)
-To evaluate the effects of various components in PTR, we conduct ablation experiments on the Chengdu and Porto datasets with sampling intervals of 2 minutes.
-We compare PTR with the following five variants.
-  • w/o explicit trajectory prompt: We remove the explicit trajectory prompt.
-  • w/o implicit trajectory prompt: We use the ‘[MASK]’ token in the BERT to represent the missing location.
-  • w/o trajectory prompt: We remove both the explicit and implicit trajectory prompt, and the missing points are represented by the ‘[MASK]’ token in BERT.
-  • w/o road network: We only use the Learnable Fourier Feature for encoding the observed trajectory points while removing the road network feature.
-  • w/o reference tokens: We remove the trajectory feature transformation layer in the trajectory embedder module.
-As shown in Table 3, each component of PTR is essential.
-The explicit trajectory prompt provides task and trajectory-related textual information, enabling PLMs to effectively understand the trajectory recovery tasks.
-Omitting this component leads to a performance decrease.
-When the implicit trajectory prompt is removed, many missing locations are replaced by the ‘[MASK]’ token in BERT, limiting the model’s ability to capture external factors for missing points, which results in suboptimal performance.
-For the trajectory point embedder, excluding road network for observed points causes the model to rely solely on latitude and longitude, neglecting the relationship between trajectories and the road network, which demonstrates that road network is crucial for accurately modeling trajectory information.
+PLMTrajRec - Randomly initialized BERT 74.11 78.62 80.83 288.4 831.9
+PLMTrajRec 81.76 84.31 88.38 187.4 366.2
+As shown in Table 5 and 8, each component of PLMTrajRec is essential.
+The IF-guided explicit trajectory prompt provides task and trajectory-related textual information, enabling PLMs to effectively understand trajectory features.
+Omitting it leads to a performance decrease.
+When the AF-guided implicit trajectory prompt is removed, many missing locations are replaced by the ‘[MASK]’ token in BERT, limiting the model’s ability to capture road conditions for missing points, which results in suboptimal performance.
+For the trajectory point embedder, excluding road network for observed points causes the model to rely solely on latitude and longitude, neglecting the relationship between trajectories and the road network, which demonstrates that road network is crucial for modeling trajectory information.
 Additionally, removing reference tokens results in a performance decline, as PLMs are unable to directly interpret trajectory data.
+Using the randomly initialized BERT in PLMTrajRec has poor performance, indicating the efficiency of the PLM for trajectory recovery.
 6. Conclusion
-In this paper, we propose PTR, a novel framework for trajectory recovery leveraging PLM.
-PTR efficiently handles trajectory recovery even with limited dense trajectory data, demonstrating strong scalability.
-The model is trained on datasets with various sparse trajectories, enabling it to effectively generalize to different sampling intervals.
-To model external factors, we introduce an implicit trajectory prompt to model road conditions, which provides rich information for missing trajectory points.
-We also develop a trajectory embedder to capture the spatial-temporal dependencies of sparse trajectories and convert them into a format interpretable by PLMs.
-Experimental results on two datasets with three different sampling intervals validate the efficacy and scalability of the proposed model.
-Appendix A Dataset
-Table 4. Dataset Description.
-Types Chengdu Porto
-Sampling interval 15s 15s
-# Trajectory 118,354 322,079
-# Road segment 2504 2224
-Latitude range 30.655 $\sim$ 30.727 41.142 $\sim$ 41.174
-Longitude range 104.043 $\sim$ 104.129 -8.652 $\sim$ -8.578
-Length / Width 8.22 km / 8.00 km 6.19 km / 3.56 km
-Appendix B Baseline Setting
-We choose the following ten methods as baselines, including five free space trajectory recovery and five map-matched trajectory recovery.
-B.1. Free-space Trajectory Recovery
+In this paper, we propose PLMTrajRec, a novel model for trajectory recovery leveraging PLM.
+PLMTrajRec effectively recovers trajectory even with limited dense trajectory data available, demonstrating strong scalability.
+The model equippes with an interval and feature-guided explicit trajectory prompt and interval-aware trajectory prompt enabling it to effectively generalize to different sampling intervals.
+Additionally, we introduce an area flow-guided implicit trajectory prompt to gather traffic flows in each region, and propose a road condition passing mechanism to infer missing-point conditions from nearby observations.
+Experimental results on two datasets with three different sampling intervals validate the effectiveness, scalability, and generalizability of the proposed model.
+Appendix A Baseline Setting
+We choose the following 12 methods as baselines, including five free space trajectory recovery and seven map-matched trajectory recovery.
+A.1. Free-space Trajectory Recovery
 Free space trajectory recovery first recovers trajectory points and then projects the trajectory onto the road network.
@@
   • AttnMove (Xia et al., 2021) + Rule uses attention to predict the missing road segments and uses the central location as the moving rate.
-B.2. Map-matched Trajectory Recovery
+A.2. Map-matched Trajectory Recovery
 Map-matched trajectory recovery can directly recover the trajectory on the road network.
@@
 It optimizes road segment and moving rate prediction through multi-task learning.
-  • T2vec (Li et al., 2018) + Decoder is a deep learning model for trajectory similarity learning.
-We use its encoder to embed the sparse trajectory and use the decoder of the MTrajRec to recover the trajectory in an autoregressive manner.
-  • T3s (Yang et al., 2021) + Decoder uses LSTM and attention mechanisms to encode sparse trajectory data effectively. Then, the MTrajRec decoder is utilized to recover the trajectory.
+  • T2vec (Li et al., 2018) is a deep learning model for trajectory similarity learning. We use its encoder to embed the sparse trajectory.
+  • T3s (Yang et al., 2021) uses LSTM and attention mechanisms to encode sparse trajectory data effectively.
+  • TERI (Chen et al., 2023a) assumes that the number of missing trajectory points is unknown and proposes a two-stage trajectory recovery framework.
+In the first stage, a transformer-based model predicts the number of points to be recovered, and in the second stage, the same framework is used for recovery trajectory coordinates.
+Here, we utilize only the second stage of TERI.
+  • TrajBERT (Si et al., 2023) employs a transformer encoder and a forward and backward neighbor selector to learn complex mobility patterns bi-directionally from sparse trajectories.
   • RNTrajRec (Chen et al., 2023b) leverages the Transformer to capture the spatial-temporal correlation of the sparse trajectories.
@@
   • MM-STGED (Wei et al., 2024) models sparse trajectories from a graph perspective and recovers trajectories by capturing micro and macro semantic information.
-Appendix C Evaluation Metrics
+Notably, despite T2vec, T3S, TERI, and TrajBERT employing different techniques to capture sparse trajectories’ spatial-temporal dependencies, they cannot directly generate the desired format for missing points.
+Therefore, after these models obtain the trajectory embeddings, we append the Decoder part of MTrajRec to output the road segment and moving ratio of the trajectory point.
+Denoted by T2v + Decoder, T3S + Decoder, TERI + Decoder, and TrajBERT + Decoder, respectively.
+Appendix B Evaluation Metrics
 We adopt five widely used metrics to evaluate the effectiveness of our model, following previous works (Wei et al., 2024; Chen et al., 2023b; Ren et al., 2021).
@@
 Lower values of these metrics indicate a higher level of accuracy in the recovery results.
-Appendix D Appendix Experiments
-D.1. Covergence Rate
-Figure 5. Convergence rate of PTR and fine-tuning for three sampling intervals on Chengdu.
-Figure 6. Convergence rate of PTR and fine-tuning for three sampling intervals on Porto.
-We evaluate the convergence speed of our model on the Porto and Chengdu datasets.
-As illustrated in Figure 5 and 6, the model reaches near-optimal performance by the 10th epoch.
-The trained model demonstrated robust accuracy across three sampling intervals (PTR: $u=1$ minute, PTR: $u=2$ minutes, and PTR: $u=4$ minutes), confirming that PTR can generalize effectively to sparse trajectories with varying sampling intervals.
-Furthermore, after fine-tuning at each sampling interval, the performance shows minimal change in performance by the 3rd epoch, further proving its fast convergence speed.
+Appendix C Example of the IF-guided Explicit Trajectory Prompt
+Consider a sparse trajectory $\mathcal{T}=\langle p_{1},\cdots,p_{N}\rangle$ of $N$ trajectory points with a sampling interval of 4 minutes, starting at 8 o’clock on Saturday and ending at 9 o’clock on Saturday.
+Our goal is to recover it within a sampling interval of 15 seconds.
+Therefore, the trajectory prompts that are related to sampling intervals are: Task Part: Sparse trajectory recovery.
+Target Part: Output the road segment and moving ratio for each point in the trajectory.
+Content Part: The sparse trajectory is sampled every four minutes and aims to recover trajectory every fifteen seconds.
+The movement feature-related trajectory prompts are: Time Part: The trajectory started at eight o’clock on Saturday and ended at nine o’clock on Saturday.
+Movement Part: Total time cost: sixty minutes zero seconds.
+Total space transfer distance: $z$ kilometers.
+Here $z={\textstyle\sum_{i=2}^{N}}\text{dist}(p_{i},p_{i-1})$, where $\mathrm{dist}(\cdot,\cdot)$ is used to calculate the distance between two trajectory points.
+Appendix D Insight about the Learnable Fourier Features
+Consider two trajectory points $x$ and $y$, and the feature mapping function $\Phi(x)=W_{\Phi}[\mathrm{cos}\text{ }xW_{r}||\mathrm{sin}\text{ }xW_{r}]$ in Learnable Fourier Features.
+The relative information $x-y$ between points $x$ and $y$ can be captured through multiplication operations, i.e.:
+(19) $\begin{split}\Phi(x)\cdot\Phi(y)&=W_{\Phi}[\mathrm{cos}\text{ }xW_{r}||\mathrm{sin}\text{ }xW_{r}]\cdot W_{\Phi}[\mathrm{cos}\text{ }yW_{r}||\mathrm{sin}\text{ }yW_{r}]\\ &=||W_{\phi}||_{2}\cdot(\mathrm{cos}\text{ }x\cdot\mathrm{cos}\text{ }y+\mathrm{sin}\text{ }x\cdot\mathrm{sin}\text{ }y)\cdot||W_{r}||_{2}\\ &=||W_{\phi}||_{2}\cdot\mathrm{cos}(x-y)\cdot||W_{r}||_{2}\end{split}$
+In our PLMTrajRec, after feature conversion, we input the trajectory feature into a pre-language training model based on BERT.
+Since there are a lot of multiplication-based attention operations in the PLM, the relative information $x-y$ can be easily modeled and utilized.
+This relative information helps infer crucial details like the distance between trajectory points, which is important for understanding movement.
+For instance, a larger distance between two points may indicate a higher likelihood of vehicle acceleration.
+Appendix E The Detailed Description of Variant
+  • w/o IF-guided explicit trajectory prompt: We remove the IF-guided explicit trajectory prompt.
+  • w/o AF-guided implicit trajectory prompt: We use the ‘[MASK]’ token in the BERT to represent the missing location.
+  • w/o dual trajectory prompts: We remove both the IF-guided explicit and implicit trajectory prompt, and the missing points are represented by the ‘[MASK]’ token in BERT.
+  • w/o road network: We only use the Learnable Fourier Feature for encoding the observed trajectory points while removing the road network feature.
+  • w/o reference tokens: We remove the trajectory feature transformation layer in the trajectory embedder module.
+  • PLMTrajRec - Randomly initialized BERT: We randomly initialize the parameter of BERT instead of using the pre-trained BERT based on large-scale corpus datasets.
+Appendix F Appendix Experiments
 Table 6. Hyperparameter range and optimal value.
@@
 Figure 7. Hyperparameter analysis of the weight $\lambda$ in the loss function.
-D.2. Hyperparameter Study
+F.1. Hyperparameter Study
 To explore the impact of hyperparameters on model performance, we conduct hyperparameter analysis using the Chengdu and Porto datasets with the sparse trajectory sampling interval of 2 minutes.
@@
 Details regarding the range of selection and the optimal values for these hyperparameters are presented in Table 6.
-D.2.1. The Number of Reference Tokens $K$
+Table 7. Scalability analysis on Chengdu dataset with 1 minute and 4 minutes sampling intervals.
+Sampling Interval Data Ratio 20% 40% 60% 80% 100%
+[…]
+$\mu=4$ minutes $\to\epsilon=15$ seconds MTrajRec 60.73 1048.0 63.58 968.5 64.73 929.4 65.39 914.6 65.79 904.4
+T3s + Decoder 60.49 1098.9 63.32 968.7 64.23 961.9 65.04 942.3 65.60 926.3
+T2vec + Decoder 62.17 966.1 64.92 946.1 64.95 972.3 65.31 946.8 66.51 915.2
+RNTrajRec 60.80 998.0 62.85 931.3 64.19 909.1 65.46 835.0 67.66 886.0
+MM-STGED 64.61 935.0 67.76 881.8 68.53 853.7 69.95 825.3 70.64 829.7
+PLMTrajRec 68.30 585.3 71.57 515.6 72.74 512.5 73.56 488.3 74.12 483.0
+$\mu=1$ minute $\to\epsilon=15$ seconds MTrajRec 75.39 937.5 78.53 835.1 80.01 794.3 80.93 725.7 81.12 718.4
+T3s + Decoder 76.49 917.1 79.08 824.9 80.65 767.8 80.82 716.5 80.90 713.0
+T2vec + Decoder 75.89 845.4 79.09 746.8 79.41 752.3 81.21 742.2 81.69 714.1
+RNTrajRec 75.65 846.2 79.48 784.8 79.61 769.3 81.74 742.5 81.88 702.5
+MM-STGED 76.02 857.7 79.86 734.7 82.25 676.4 83.44 663.0 84.26 633.5
+PLMTrajRec 81.62 397.5 84.59 352.0 85.75 328.7 86.73 305.4 87.17 290.8
+F.1.1. The Number of Reference Tokens $K$
 We set the parameter $K$ from the set $\{128,256,512,1024\}$ to explore its impact on trajectory recovery.
@@
 This suggests that having a larger token space aids in accurately representing trajectory features.
-When $K$ is larger than 512, the performance improvement of PTR becomes marginal, while the computational cost will increase.
+When $K$ is larger than 512, the performance improvement of PLMTrajRec becomes marginal, while the computational cost will increase.
 To balance the performance and efficiency of the model, we set $K$ to 512.
-D.2.2. The Rank $r$ of LoRA
+F.1.2. The Rank $r$ of LoRA
 As shown in Figure 6, we set the range of $r$ to 4, 8, 16, and 64 to explore its impact.
@@
 Therefore, we set $r=8$ to balance the performance and resources.
-D.2.3. The Loss Weight $\lambda$
+F.1.3. The Loss Weight $\lambda$
 As shown in Figure 7, as the loss function weight $\lambda$ increases, the model performance first improves and then decreases.
@@
 To balance these two tasks, we set the value of $\lambda$ to 10.
-Figure 10. Case study on the Chengdu dataset. Red points represent the truth trajectory points and blue points represent the recovered trajectory points.
-D.3. Case Study
-We conduct a case study on the Chengdu dataset to visualize the trajectory recovery performance with various baselines.
-As shown in Figure 10, we draw the truth and recovered trajectory, where red points represent the truth trajectory points and blue points indicate the recovered trajectory points.
-We find that the recovered trajectory aligns well with the road network, demonstrating the effectiveness of using road segment and moving rate to represent trajectory point.
-To facilitate a more intuitive comparison of recovery performance among different models, we focus on two distinct regions, labeled as A and B, for visual analysis.
-In region A, characterized by a relatively simple road network structure, all models can accurately recover road segments.
-Among them, PTR maps trajectory points better by leveraging PLMs.
-In contrast, region B exhibits a more intricate road network with multiple accessible routes.
-RNTrajRec captures spatial-temporal correlations of trajectories that are not adequate and recovers incorrect road segments.
-PTR not only excels in road segment recovery but also in accurately matching the actual trajectory points.
-This success can be attributed to its ability to model missing trajectory points through the implicit trajectory prompt, thereby introducing more valuable information and improving performance.
-D.4. Computational Cost
-We use an NVIDIA RTX A4000 card to conduct cost analysis on the Chengdu and Porto datasets with a 2-minute sampling interval by considering model size, training time, and inference time.
-As shown in Table 6, compared to the end-to-end deep learning model, PTR has the largest number of parameters due to its utilization of a pre-trained language model based on the BERT-small framework.
-Despite having a greater number of parameters, PTR exhibits quicker training and testing times as it can output results for all trajectory points at once, and the majority of parameters are frozen, unlike other models that necessitate autoregressive generation.
-It significantly accelerates both model training and inference processes.
+F.2. Scalability Study on 1-minute and 4-minute Sampling Interval
+The result of the scalability study on the sparse trajectory with 1-minute and 4 minutes, as shown in Table 7, the conclusion is consistent with the 2-minute sampling interval.
+The experiment results prove the scalability of PLMTrajRec.
+Table 8. Ablation study on Porto dataset with sampling intervals at 2 minutes.
+[…]
+w/o IF-guided explicit trajectory prompt 66.13 66.19 81.98 157.9 314.6
+w/o AF-guided implicit trajectory prompt 66.18 66.22 82.01 149.6 298.1
+w/o dual trajectory prompts 65.52 65.89 81.51 171.0 339.4
+w/o road network 66.11 66.27 81.90 150.7 302.8
+w/o reference tokens 65.89 65.07 80.57 165.4 327.7
+PLMTrajRec - Randomly initialized BERT 62.06 64.18 79.26 179.3 413.9
+PLMTrajRec 66.40 66.52 82.14 141.9 294.6
+F.3. Ablation Study on the Porto
+The result of the ablation study on the Porto dataset, as shown in Table 8, the conclusion is consistent with the Chengdu dataset.
+Each component contributes to PLMTrajRec’s performance, and removing any component results in a decline in model effectiveness.
 Table 9. Computation Cost.
@@
 MM-STGED 34 / 31 41.27 / 138.33 16.27 / 30.58
-PTR 176 / 175 24.18 / 54.14 2.56 / 10.38
+PLMTrajRec 176 / 175 24.18 / 54.14 2.56 / 10.38
+F.4. Computational Cost
+We use an NVIDIA RTX A4000 card to conduct cost analysis on the Chengdu and Porto datasets with a 2-minute sampling interval by considering model size, training time, and inference time.
+As shown in Table 9, compared to the end-to-end deep learning model, PLMTrajRec has the largest number of parameters due to its utilization of a pre-trained language model based on the BERT-small framework.
+Despite having a greater number of parameters, PLMTrajRec exhibits quicker training and testing times as it can output results for all trajectory points at once, and the majority of parameters are frozen, unlike other models that necessitate autoregressive generation.
+It significantly accelerates both model training and inference processes.
+Figure 8. Case study on the Chengdu dataset. Red points represent the truth trajectory points and blue points represent the recovered trajectory points.
+F.5. Case Study
+We conduct a case study on the Chengdu dataset to visualize the trajectory recovery performance with various baselines.
+As shown in Figure 8, we draw the truth and recovered trajectory, where red points represent the truth trajectory points and blue points indicate the recovered trajectory points.
+We find that the recovered trajectory aligns well with the road network, demonstrating the effectiveness of using road segment and moving rate to represent trajectory point.
+To facilitate a more intuitive comparison of recovery performance among different models, we focus on two distinct regions, labeled as A and B, for visual analysis.
+In region A, characterized by a relatively simple road network structure, all models can accurately recover road segments.
+Among them, PLMTrajRec maps trajectory points better by leveraging PLMs.
+In contrast, region B exhibits a more intricate road network with multiple accessible routes.
+RNTrajRec captures spatial-temporal correlations of trajectories that are not adequate and recovers incorrect road segments.
+PLMTrajRec not only excels in road segment recovery but also in accurately matching the actual trajectory points.
+This success can be attributed to its ability to model missing trajectory points through the implicit trajectory prompt, thereby introducing more valuable information and improving performance.